### v1

# MOTGNN: Interpretable Graph Neural Networks for Multi-Omics Disease Classification

Abstract.
Integrating multi-omics data, such as DNA methylation, mRNA expression, and microRNA (miRNA) expression, offers a comprehensive view of the biological mechanisms underlying disease. However, the high dimensionality and complex interactions among omics layers present major challenges for predictive modeling. We propose Multi-Omics integration with Tree-generated Graph Neural Network (MOTGNN), a novel and interpretable framework for binary disease classification. MOTGNN employs eXtreme Gradient Boosting (XGBoost) to perform omics-specific supervised graph construction, followed by modality-specific Graph Neural Networks (GNNs) for hierarchical representation learning, and a deep feedforward network for cross-omics integration. On three real-world disease datasets, MOTGNN outperforms state-of-the-art baselines by 5-10% in accuracy, ROC-AUC, and F1-score, and remains robust to severe class imbalance (e.g., 87.2% vs. 33.4% F1 on imbalanced data). The model maintains computational efficiency through sparse graphs (2.1-2.8 edges per node) and provides built-in interpretability, revealing both top-ranked biomarkers and the relative contributions of each omics modality. These results highlight MOTGNN’s potential to improve both predictive accuracy and interpretability in multi-omics disease modeling.
Graph neural networks, Multi-omics integration, XGBoost, Disease classification, Model interpretability
conference: ; ;
1. Introduction
A comprehensive understanding of complex diseases, such as cancer, cardiovascular disease, and neurodegenerative disorders, is critical for advancing public health and enabling precision medicine. These diseases are not driven by isolated molecular events but emerge from intricate interactions among genomic, epigenomic, and transcriptomic factors (Subramanian et al., 2020). Recent advances in high-throughput biotechnologies have enabled the simultaneous collection of different layers of molecular regulation, such as gene expression (messenger RNA), epigenetic modifications (DNA methylation), and non-coding RNA activity (e.g., microRNA), from the same set of biological samples (Hood and Rowen, 2013). These omics datasets capture intricate relationships among biological entities and molecular processes, offering insights into disease mechanisms. While individual omics types provide valuable information, different modalities reflect distinct biological processes and contribute complementary information. Integrating multiple omics (multi-omics) data is essential to comprehensively understand the complex, multi-layered systems underlying disease processes and their progression (Waqas et al., 2024). However, integrating heterogeneous omics data presents unique challenges. A fundamental issue is that the number of samples ($n$) is substantially smaller than the number of features ($p$). For example, a study may involve only a few hundred participants while measuring tens of thousands of molecular features. This high dimensionality complicates learning and increases the risk of overfitting. Moreover, not all measured features are biologically relevant; noisy or redundant features can obscure meaningful patterns. Although prior biological knowledge, such as gene-gene or protein-protein interactions, has been identified experimentally, our knowledge is constrained by experimental design and equipment constraints. Consequently, there is a growing need for computational models that can reliably infer biologically meaningful structures and interactions from multi-omics data (Jiang et al., 2025).
Traditional machine learning and deep learning models have been extensively applied to address the high dimensionality and complexity of omics data (Ayman et al., 2023; Zaghlool and Attallah, 2022), but most of these approaches assume a Euclidean data structure. Classical machine learning methods, such as random forests (Breiman, 2001) and gradient boosting methods, such as XGBoost (Chen and Guestrin, 2016), operate by constructing ensembles of decision trees and are effective for classification tasks, but they typically treat features as independent and fail to capture structural relationships among them. Deep learning models (Lecun et al., 2015), such as deep feedforward networks (DFNs) (Goodfellow et al., 2016), convolutional neural networks, and recurrent neural networks, can model nonlinear feature dependencies but are not designed to process non-Euclidean data structures such as graphs. These approaches typically require large training sets to avoid overfitting. Moreover, their black-box nature often limits interpretability, which is essential in biomedical applications where transparency and reproducibility are critical. Furthermore, concatenating features from different omics modalities as input can obscure modality-specific relationships and introduce additional noise. Graph-based approaches offer a natural alternative for modeling complex biological systems. In omics applications, nodes can represent features (e.g., genes) or samples, and edges encode various relationships, such as similarity, co-expression, or learned interactions. Graph Neural Networks (GNNs) (Bronstein et al., 2017; Wu et al., 2020; Zhang et al., 2020; Zhou et al., 2020; Chen et al., 2023) have emerged as powerful tools for learning from graph-structured data by recursively aggregating information from each node’s neighborhood. GNNs capture both local and global graph topology and have found increasing use in biomedical applications, including molecular modeling, medical imaging, drug discovery, and disease prediction (Zhang et al., 2021; Johnson et al., 2024; Waqas et al., 2024; Zhong et al., 2025). Foundational models such as message passing neural networks (MPNNs) (Gilmer et al., 2017), graph convolutional networks (GCNs) (Kipf and Welling, 2017), graph attention networks (GATs) (Petar et al., 2018), GraphSAGE (Hamilton et al., 2017), and graph isomorphism networks (GINs) (Xu et al., 2019) have significantly advanced the expressiveness and scalability of GNN-based learning.
Despite recent advances, applying GNNs to multi-omics data remains challenging and relatively underexplored (Zhang et al., 2021; Valous et al., 2024; Paul et al., 2024; Ballard et al., 2024). The irregular nature of graph-structured data further complicates this task, as each node may have a variable number of neighbors, and many standard deep learning operations are not directly applicable in this setting (Chen et al., 2023). Although several GNN-based frameworks have been proposed for multi-omics integration, many of these works exhibit substantial limitations, which we summarize into three main areas: (1) Inadequate modeling of high-dimensional, heterogeneous omics data: To address high dimensionality (i.e., the “small $n$, large pp” issue), existing methods typically apply simple filtering techniques, such as removing low-variance or zero-valued features, or selecting features based on univariate statistical tests. Some approaches flatten multi-omics data by concatenating omics layers before graph construction, which weakens modality-specific signals and ignores biological heterogeneity. Others rely heavily on hand-crafted similarity graphs (e.g., Pearson or cosine similarity), which may miss nonlinear, task-specific relationships, ignore the sparse and scale-free nature of biological graphs, and are sensitive to arbitrary thresholding or preprocessing pipelines. (2) Vulnerability to class imbalance: While class imbalance is prevalent in biomedical datasets (e.g., rare diseases, early-stage cancer detections, or minority subtypes), it is rarely addressed during model development or evaluation. Many existing models optimize for overall accuracy, which can be misleading when predictions favor the majority class. In these scenarios, even high accuracy may reflect poor minority class detection. As a result, these approaches often lack robustness and generalizability in imbalanced settings, where accurate prediction of minority cases is of greater biological and clinical importance. (3) Limited interpretability and biological insight: Interpretability is essential in biomedical research, as understanding which features or omics modalities drive predictions is essential for trust, transparency, and downstream biological discovery. However, most existing GNN-based models lack built-in interpretability. Instead, they rely on computationally intensive post hoc analyses or ablation studies to infer importance, which are not scalable and may introduce instability. In addition, they typically treat all omics types equally, failing to uncover their relative contributions. To address these limitations, we propose MOTGNN, a novel and interpretable graph-based framework for multi-omics data integration that combines supervised graph construction with hierarchical representation learning. Our key contributions are as follows:
  • Modality-specific modeling with supervised graph learning: We utilize XGBoost to filter informative features for each omics type, and leverage the structure of the trained decision trees to construct sparse, supervised, and modality-specific graphs. This approach preserves biologically meaningful interactions while reducing noise and redundancy. We then apply GNNs to learn latent representations from each graph, which are integrated through a DFN to model cross-omics interactions.
  • Imbalance-robust architecture: MOTGNN consistently outperforms existing models across all datasets, achieving 10-50% improvement in F1-score on class-imbalanced real-world data. Its design mitigates overfitting to the dominant classes and remains effective in detecting minority class signals, addressing a critical limitation in biomedical classification tasks.
  • Integrated interpretability at feature and omics levels: MOTGNN enables end-to-end interpretability without requiring post hoc analysis. It provides (i) feature-level importance scores to identify top-ranked biomarkers within each modality, and (ii) omics-level contribution scores, offering insights into which data types (e.g., methylation, mRNA, or miRNA) are most influential for prediction. This layered interpretability potentially supports downstream biological discovery and facilitates clinical relevance.
Figure 1. Overview of the proposed MOTGNN framework, which involves three key modules: (i) XGBoost for graph construction within each omics modality; (ii) GNN for learning modality-specific embeddings by encoding both feature graphs and associated data matrices; and (iii) DFN that integrates the learned embeddings and performs classification.
[…]
[MOTGNN architecture]The MOTGNN framework includes XGBoost for graph construction, a GNN module for encoding, and a DFN for integration and classification.
2. Related Work
Many advanced deep learning methods have been developed by incorporating novel model components to analyze omics data. For example, MOCAT (Yao et al., 2024) leverages an autoencoder, three omics-specific classifiers, a multi-head attention mechanism, and a confidence network (ConfNet) to integrate and learn from multi-omics data (e.g., methylation, mRNA, and miRNA). However, it lacks the ability to capture topological information and thus may miss important structural insights. Some GNN-based methods have been developed for single-omics data analysis in disease classification or cancer subtype prediction. For example, forgenet (Kong and Yu, 2020) constructs a supervised directed graph using a tree-ensemble model, but it uses the original high-dimensional data without dimension reduction, leading to high computational costs. Moreover, it only uses a single omics type, which provides limited information compared to multi-omics integration. DMGCN (Yang et al., 2021) integrates multiple biological networks via GCNs, but also focuses on a single omics modality and does not utilize cross-omics information.
More recent GNN-based methods aim to integrate multi-omics data for disease classification, cancer subtype analysis, and survival prediction. MOGONET (Wang et al., 2021) proposes a GCN-based framework to integrate multi-omics data (methylation, mRNA, and miRNA), using a cross-omics discovery tensor and a view correlation discovery network for final label prediction. However, it identifies important features through computationally expensive ablation studies. In addition, its graph construction is based on cosine similarity with a manually selected threshold determined by a hyperparameter (edge-to-node ratio), which must be tuned. MODILM (Zhong et al., 2023) shares a similar architecture with MOGONET but uses GAT instead of GCN. CLCLSA (Zhao et al., 2024) handles incomplete multi-omics classification using cross-omics autoencoders to recover missing omics representations, contrastive learning to align representations, and self-attention at both the feature and omics levels for feature selection, but does not explicitly leverage graph structures. SUPREME (Kesimoglu and Bozdag, 2023) employs GCNs and patient similarity networks, integrating seven types of omics data for cancer subtype prediction and survival analysis. However, it also relies on ablation studies to determine the most effective omics combinations. MOGAT (Tanvir et al., 2024) focuses on survival analysis with a similar structure to SUPREME, but uses GAT instead of GCN. It conducted integration studies with two to eight omics types and found that using three omics achieved the best performance for breast cancer subtype prediction, suggesting that adding more omics data does not always yield better results. However, MOGAT does not address model interpretability, as it does not identify important biomarkers. DeepMoIC (Wu et al., 2024) integrates three types of omics data to address cancer subtype classification by combining autoencoders, similarity network fusion, and a deep GCN with residual connections and identity mappings. However, it constructs patient similarity networks to guide the learning and lacks omics-wise interpretability. MoGCN (Li et al., 2022) targets breast cancer subtype classification using multi-omics data (copy number variation, RPPA, and RNA-seq). It applies autoencoders for dimensionality reduction and similarity network fusion for graph construction, and trains a GCN model on the resulting fused network. However, its graph construction still relies on distance-based methods.
While these models demonstrate promising results, several common limitations remain: They require extensive data preprocessing, and the quality of such preprocessing can significantly affect downstream model performance. Their graph construction approaches rely on simple similarity metrics, which often fail to capture task-relevant nonlinear interactions and can introduce bias due to arbitrary thresholding. They rarely address the class imbalance issue, which is prevalent in biomedical datasets and can significantly impact performance. They commonly depend on ablation studies to identify the optimal combination of omics data types, a computationally expensive process that is not easily scalable.
3. Methodology
We summarized the key notations and their associated descriptions in Table 1. A graph is denoted as $G(V,E)$, where $V$ represents the set of vertices (or nodes) and $E$ denotes the set of edges. Graph structure is encoded by its adjacency matrix $A$, in which the $(i,j)$-th element is one if node $i$ is connected to node $j$, and zero otherwise. For undirected graphs, $A$ is symmetric. The adjacency matrix augmented with self-loops is denoted as $\tilde{A}$, where all diagonal elements are set to one to include self-connections.
Table 1. Key notations and descriptions
Notation Description
$G$ or $G(V,E)$ Graph
$V$ Set of vertices/nodes in $G$
$E$ Set of edges in $G$
AA Adjacency matrix of $G$
$\tilde{A}$ Adjacency matrix with self-loops
$X_{i}\in\mathbb{R}^{n\times p_{i}}$ Omics data matrix for modality $i=1,2,3$
$n$ Number of samples (rows in $X_{i}$)
$p_{i}$ Number of features in $X_{i}$
$Y$ Binary outcome variable
$X_{i}^{*}\in\mathbb{R}^{n\times p_{i}^{*}}$ Dimension-reduced representation of $X_{i}$
$p_{i}^{*}$ Reduced feature dimension for $X_{i}$
$G_{i}$ or $G_{i}(V_{i},E_{i})$ Graph constructed for $X_{i}^{*}$
$V_{i}$ Set of vertices/nodes in $G_{i}$
$E_{i}$ Set of edges in $G_{i}$
$Z_{i}$ Embedding learned from $G_{i}$ via GNN
$Z$ Concatenated representation of all $Z_{i}$
$m_{i}$ Edge-to-node ratio, i.e., $m_{i}=\frac{E_{i}}{p_{i}^{*}}$
$RIG_{i}$ Relative importance of graph $G_{i}$
$\odot$ Element-wise (Hadamard) product
$I(\cdot)$ Indicator function
3.1. Proposed Model
We propose a Multi-Omics integration framework with Tree-generated Graph Neural Network (MOTGNN) model for binary disease classification. MOTGNN first leverages eXtreme Gradient Boosting (XGBoost) to select informative features in a supervised manner, as the tree-based feature selection captures strong predictive signals. It then utilizes GNNs to learn omics-specific representations from graphs constructed using the structure of trained XGBoost trees. This combination allows the model to extract both local and global patterns from each omics layer while maintaining biological interpretability.
The framework of MOTGNN is shown in Figure 1, which consists of three primary modules: (1) An XGBoost module for feature graph construction; (2) A GNN module to encode each omics-specific graph and its associated data matrix; and (3) A DFN module to integrate and classify the learned embeddings. This modular architecture allows MOTGNN to handle high-dimensional, heterogeneous multi-omics data efficiently while providing interpretable insights at both the feature and omics-layer levels. Algorithm 1 demonstrates the steps of the overall structure of MOTGNN.
Algorithm 1 MOTGNN
1:Omics datasets: $X_{1}\in\mathbb{R}^{n\times p_{1}}$, $X_{2}\in\mathbb{R}^{n\times p_{2}}$, $X_{3}\in\mathbb{R}^{n\times p_{3}}$; Binary disease labels: $Y\in\mathbb{R}^{n\times 1}$
2:Predicted labels $\hat{Y}$
3:Train three XGBoost models using $X_{1}$, $X_{2}$, and $X_{3}$ to predict $Y$
4:Construct omic-specific feature graphs: $G_{1}(V_{1},E_{1})$, $G_{2}(V_{2},E_{2})$, and $G_{3}(V_{3},E_{3})$ by taking the union of decision paths from all trees trained in Step 1, along with the corresponding dimension-reduced datasets: $X_{1}^{*}\in\mathbb{R}^{n\times p_{1}^{*}}$, $X_{2}^{*}\in\mathbb{R}^{n\times p_{2}^{*}}$, and $X_{3}^{*}\in\mathbb{R}^{n\times p_{3}^{*}}$
5:for each $(X_{i}^{*},G_{i})$, $i=1,2,3$ do
6:  Feed $X_{i}^{*}$ and $\tilde{A}_{i}$ (augmented adjacency matrix of $G_{i}$ with self-loops) into a GNN
7:  Obtain learned embedding $Z_{i}$
8:end for
9:Concatenate $Z_{1}$, $Z_{2}$, and $Z_{3}$ to form a unified representation
10:Pass the concatenated representation $Z$ into a DFN to perform final binary classification
11:return Predicted classification labels $\hat{Y}$
3.2. Graph Construction
XGBoost (Chen and Guestrin, 2016) is a powerful ensemble method based on gradient-boosted decision trees. Compared to traditional gradient boosting machines (GBMs) (Friedman, 2002), XGBoost introduces a regularized learning objective to reduce overfitting and improve generalization. In MOTGNN, we employ three XGBoost models, each trained independently to perform binary classification on the omics-specific datasets $X_{1}$, $X_{2}$, and $X_{3}$ alongside the binary response $Y$. Unlike random forest, which generates independent trees, each trained XGBoost model generates $M$ sequential decision trees, where every tree improves upon the residuals of the previous ones. During training, each tree selects a subset of features for node splits. We collect the union of all features used across the $M$ trees and define these as the selected features. This yields a supervised, tree-guided feature selection process that reduces the original feature dimensions from $p_{1}$, $p_{2}$, and $p_{3}$ to $p_{1}^{*}$, $p_{2}^{*}$, and $p_{3}^{*}$, respectively, where $p_{i}^{*}<p_{i}$ and $p_{1}^{*}\neq p_{2}^{*}\neq p_{3}^{*}$. The reduced dimensions preserve features with higher predictive importance, facilitating downstream graph construction and neural modeling.
Inspired by forgeNet (Kong and Yu, 2020), which constructs feature graphs based on the structure of ensemble trees, we adapt this approach to build omics-specific undirected graphs from XGBoost models. Specifically, for each omics dataset, we treat each trained decision tree as a graph, where features involved in splits are nodes, and parent-child relationships form the undirected edges. Taking $X_{1}\in\mathbb{R}^{n\times p_{1}}$ along with the binary outcome $Y$ as an example, suppose the XGBoost model generates $M$ sequential decision trees. Each tree is considered an undirected graph $G^{m}=(V^{m},E^{m})$ for $m=1,2,\ldots,M$. We then take the union of all tree-level graphs to form the aggregated graph $G_{1}=(V_{1},E_{1})$ as:
$G_{1}(V_{1},E_{1})=G_{1}(\bigcup_{m=1}^{M}V^{m},\bigcup_{m=1}^{M}E^{m}),$
where the number of nodes in $G_{1}$ is $|V_{1}|=p_{1}^{*}$, and $p_{1}^{*}<p_{1}$. We also add self-loops to the adjacency matrices to preserve self-information in GNN message passing. A similar procedure generates $G_{2}$ and $G_{3}$ for the remaining two omics datasets.
3.3. Graph Neural Network and Feature Fusion
We utilize the graph-embedded deep feedforward network (GEDFN) model (Kong and Yu, 2018) as the core GNN component used in MOTGNN. Unlike standard fully connected architectures, GEDFN directly incorporates the adjacency matrix into the weight connections between the input and the first hidden layers. This design enables the model to respect the feature-level relationships encoded in the graph while enforcing the sparsity, a common assumption for omics-based feature graphs. The number of neurons in the first hidden layer must be the same as the input dimension. Formally, given the reduced input data $X^{*}\in\mathbb{R}^{n\times p^{*}}$ and its graph adjacency matrix $\tilde{A}\in\mathbb{R}^{p^{*}\times p^{*}}$ (augmented with self-loops), the first hidden layer $Z_{1}\in\mathbb{R}^{n\times p^{*}}$ is computed as
$Z_{1}=\sigma(X^{*}(W_{in}\odot\tilde{A})+b_{in}),$
where $W_{in}\in\mathbb{R}^{p^{*}\times p^{*}}$ and $b_{in}\in\mathbb{R}^{n\times p^{*}}$ are the weight and bias matrices, $\odot$ denotes element-wise (Hadamard) multiplication, and $\sigma(\cdot)$ is the activation function. The design of GEDFN shares conceptual similarities with the graph attention network (GAT) formulation discussed in (Chen et al., 2023), where the GAT layer is represented as
$Z_{1}=\sigma((W_{attention}\odot A)X+b_{in}),$
with $W_{attention}\in\mathbb{R}^{n\times n}$ and the nodes representing samples rather than features. In this formulation, GAT dynamically learns attention weights over neighbors. Similarly, GEDFN enables dynamic weight learning while ensuring that connections between features and the first hidden layer are strictly supported by the feature graph. This yields a sparse and interpretable architecture tailored to omics data.
Following the training of three omics-specific GNNs, each on its respective dimension-reduced input and constructed graph, we obtain three last hidden representations: $Z_{1}$, $Z_{2}$, and $Z_{3}$. These embeddings are concatenated to form a unified representation:
$Z=[Z_{1}|Z_{2}|Z_{3}],$
where $Z$ denotes the concatenated neurons, and is then passed to a DFN, which consists of fully connected layers and a final softmax output to perform binary disease classification.
Table 2. Data description of three TCGA cancer datasets (COADREAD, LGG, and STAD) regarding sample sizes, class distributions, the number of original features, and the number of preprocessed features for each omics type (methylation, mRNA, and miRNA).
Abbr. Data Name Samples Class Balance Original Features Preprocessed Features
(0:1) (meth:mRNA:miRNA) (meth:mRNA:miRNA)
COADREAD Colorectal and Rectal Adenocarcinoma 332 COAD:READ = 254:78 20,113:20,531:420 2,000:2,000:420
LGG Low-Grade Glioma 524 Grade 2:Grade 3 = 255:269 20,531:20,114:548 2,000:2,000:548
STAD Stomach Adenocarcinoma 371 ADC:IAC = 205:166 20,101:20,531:507 2,000:2,000:507
Note: meth = DNA methylation, mRNA = gene expression, miRNA = microRNA expression. The processed data were obtained from the Broad GDAC Firehose (Lu et al., 2023).
Figure 2. The histogram of methylation, mRNA, and miRNA on preprocessed COADREAD, LGG, and STAD datasets. Each omics dataset was individually scaled to [0, 1]. Different omics types demonstrate distinct distribution patterns.
[…]
[Omics distribution]The histograms of methylation, mRNA, and miRNA on COADREAD, LGG, and STAD datasets. Each omics dataset was individually scaled to the [0, 1] range using min-max normalization. Different omics data demonstrate distinct distribution patterns.
3.4. Interpretability on Features and Graphs
To interpret model predictions and evaluate the contribution of features and graphs, we adopt and extend the connection weights method (Olden and Jackson, 2002). Feature importance is assessed by summing the absolute weights connecting each input neuron to the first hidden layer, while accounting for adjacency-based sparsity. For feature $j$ in $X_{i}$, the importance score $IF_{j}^{i}$ is defined as:
$IF_{j}^{i}=\sum_{u=1}^{p_{i}}\left|W_{ju}^{(in)}I(\tilde{A}_{ju}^{i}=1)\right|,$
where $W^{(\text{in})}$ is the input weight matrix and $I(\cdot)$ is the indicator function. After training the model, each input feature is assigned an importance score. These scores can be used to rank the features, allowing the selection of top biomarkers for further biological investigation. To evaluate the relative importance of the graph for each omics modality, we compute the $L_{1}$ norm of the weight matrix connecting the GNN’s last hidden representations ($Z_{i}$) to the DFN layer. The relative graph importance (RIG) is then:
$RIG_{i}=\frac{\|W_{Z_{i}\leftrightarrow f}\|_{1}}{\sum_{i=1}^{3}\|W_{Z_{i}\leftrightarrow f}\|_{1}},$
where $i=1,2,3$, $W_{Z_{i}\leftrightarrow f}$ is the weight matrix connecting $Z_{i}$ and DFN, and $\|\cdot\|_{1}$ is the $L_{1}$ norm that sums the absolute values of its inside matrix. The total contribution sums to one across all three graphs.
3.5. Technical Details
We implement the xgboost Python library (Chen and Guestrin, 2016) for the XGBoost module in MOTGNN. We tuned the number of estimators (determine trees). We employ the Tensorflow library (Abadi et al., 2016) for training GNN and DFN modules. Model optimization uses the Adam optimizer (Kingma and Ba, 2014) with mini-batch training. The nonlinear activation function uses the Rectified Linear Unit (ReLU) function: $\sigma(x)=max(0,x)$ (Nair and Hinton, 2010). For classification, the final output layer uses a softmax function, and the loss is computed using binary cross-entropy:
$L(Y,\hat{Y})=-\frac{1}{n}\sum_{i=1}^{n}\left\{y_{i}\ln(\hat{y}_{i})+(1-y_{i})\ln(1-\hat{y}_{i})\right\}.$
We tune hyperparameters across three main categories: (1) architectural (number of layers and hidden neurons), (2) training-related (learning rate, batch size, and epochs), and (3) regularization-related (dropout and L2 penalty). To avoid overfitting, we apply strategies including dropout, batch normalization, L2 regularization, and early stopping. Data are split into training, validation, and test sets in a 60%:20%:20% ratio, with stratified sampling to address class imbalance. We have conducted 20 independent and reproducible train-validation-test splits. Hyperparameters are tuned via grid search in a reasonable space and based on experience.
4. Real Data Applications
4.1. Data Description
We evaluate our method on three real-world cancer datasets from The Cancer Genome Atlas (TCGA) (Weinstein et al., 2013): COADREAD, STAD, and LGG. The processed multi-omics data and label annotations were obtained from the Broad GDAC Firehose (https://gdac.broadinstitute.org/). Each dataset includes matched DNA methylation ($X_{1}$), mRNA expression ($X_{2}$), and miRNA expression ($X_{3}$) data. Only samples with complete data across all three omics types were included in our analysis. The COADREAD dataset combines COlorectal ADenocarcinoma (COAD) and REctal ADenocarcinoma (READ) into a binary classification task, with 254 COAD and 78 READ samples. The LGG dataset involves grade classification for low-grade gliomas, with 255 Grade 2 and 269 Grade 3 samples. The STAD dataset targets the classification of STomach ADenocarcinoma subtypes, with 205 samples from adenocarcinoma (ADC) and 166 from intestinal adenocarcinoma (IAC). Among the three datasets, COADREAD has the most imbalanced class distribution.
Table 2 summarizes the datasets regarding sample sizes, class distributions, the number of original features, and the number of preprocessed features for each omics type. Data preprocessing includes quality control steps to remove noisy and redundant features, such as filtering out zero-expression signals, low-variance features, and handling multiple comparisons (Lu et al., 2023). Each omics dataset was independently normalized to the range [0, 1] using min-max scaling. The preprocessed multi-omics data shared in (Lu et al., 2023) contains minor inconsistencies in sample sizes across omics modalities. For the COADREAD dataset, meth and mRNA data include 332 samples, while miRNA contains 337 samples. We addressed this discrepancy by removing the five extra miRNA samples to ensure alignment. Similarly, in the STAD dataset, meth and mRNA data comprise 371 samples, whereas miRNA includes 372 samples. Here, we excluded one redundant miRNA sample to maintain consistency.
The histograms in Figure 2 show the distribution of feature values across three omics types (methylation, mRNA, and miRNA) for the COADREAD, LGG, and STAD datasets. The distinct distributional patterns observed across omics types within each dataset highlight each modality’s inherent heterogeneity and diverse data characteristics. These differences across the three datasets help evaluate the potential of MOTGNN’s generalizability.
4.2. Comparison and Evaluation Metrics
We evaluate the performance of our proposed model, MOTGNN, in comparison to several baseline machine learning and deep learning models: XGBoost (XGB) (Chen and Guestrin, 2016), random forest (RF) (Breiman, 2001), deep feedforward network (DFN) (Goodfellow et al., 2016), and graph convolutional network (GCN) (Kipf and Welling, 2017). The number of trees for XGBoost and RF is both at the default, which is 100. The number of hidden layers for the DFN and GCN models is two, with several hidden neurons of 64 per layer. All baseline models input the column-wise concatenation of the three preprocessed omics data types (methylation, mRNA, and miRNA). Table 3 summarizes the tuning ranges and selected values (optimal choices in bold). We report three standard evaluation metrics: accuracy, ROC-AUC, and F1-score. Accuracy measures the proportion of correctly classified samples. ROC-AUC (area under the receiver operating characteristic curve) (Green et al., 1966) quantifies the model’s ability to distinguish between classes. The F1-score is defined as $F1=\frac{2\times\text{precision}\times\text{recall}}{\text{precision}+\text{recall}},$ where precision refers to the positive predictive value, and recall corresponds to sensitivity or true positive rate. As the harmonic mean of precision and recall, the F1-score provides a balanced evaluation, especially in scenarios with imbalanced class distributions, and mitigates bias from simply predicting the majority class. These metrics comprehensively assess classification performance across different datasets and modeling approaches. We conducted all computations on the Falcon supercomputer (Idaho C3+3 Collaboration, 2022), requesting four CPUs and 20 GB of memory per job.
Table 3. Hyperparameter tuning for XGBoost and neural networks components in MOTGNN on COADREAD, LGG, and STAD datasets. The optimal choices are in bold.
Hyperparameter Tuning Range
XGBoost Parameters
Number of Trees 100, 1000
Neural Network Architecture
Depth (hidden layers) 1, 2, 3
Width (neurons per layer) 32, 64, 128
Activation Function ReLU, leakyReLU
Training Parameters
Learning Rate 0.001, 0.0005, 0.0001
Batch Size 8, 16, 32
Training Epochs 200, 300, 500
Regularization and Early Stopping
Dropout Rate 0.2, 0.3, 0.5
L2 Regularization Lambda 0.01
Early Stopping Patience 5, 10
Early Stopping Min Delta 0.001, 0.005
Figure 3. Feature dimensions before and after XGBoost on COADREAD, LGG, and STAD datasets. The bar plots compare the preprocessed feature dimensions (pre-selection) against the reduced dimensions (post-selection).
[…]
[Dimension reduction]Feature Dimension before and after XGBoost selection
5. Results and Discussion
Table 4. Structural properties of sparse graphs constructed by MOTGNN for COADREAD, LGG, and STAD datasets.
Dataset Edges Edge/Node
($E_{1}$$E_{2}$$E_{3}$) ($m_{1}$$m_{2}$$m_{3}$)
COADREAD 260:318:190 2.13:2.11:2.35
LGG 880:849:856 2.26:2.18:2.79
STAD 674:646:639 2.17:2.10:2.61
Note: The number of edges includes the self-loops, $E_{i}$ is the set of edges for Graph $G_{i}$, $m_{i}$ is the edge to node ratio with $m_{i}=\frac{E_{i}}{p_{i}^{*}}$.
Figure 4. Relative graph importance of different omics types (methylation, mRNA, and miRNA) in MOTGNN’s classification performance across COADREAD, LGG, and STAD datasets. Bar heights represent each omics layer’s contribution, revealing dataset-specific biological dominance.
[…]
[Relative graph importance]Relative Graph Importance of different omics type (meth, mRNA, and miRNA) for Classification Task on COADREAD, LGG, and STAD datasets
Note: $G_{i}$ = Graphs from reduced features (methylation:mRNA:miRNA), $RIG_{i}$ = Relative importance of graph $G_{i}$ in MOTGNN (normalized to sum to 1.0 per dataset).
Table 5. Classification performance comparison on COADREAD, LGG, and STAD datasets across different evaluation metrics (Mean ± SD [95% CI]). MOTGNN consistently outperforms baseline models across all datasets, highlighting its robustness in integrating multi-omics data (Best results are bolded)
Dataset Model Accuracy ROC-AUC F1-Score
COADREAD XGBoost 0.910 ± 0.034 [0.894, 0.926] 0.964 ± 0.022 [0.953, 0.975] 0.779 ± 0.094 [0.735, 0.823]
RF 0.801 ± 0.025 [0.789, 0.813] 0.891 ± 0.038 [0.873, 0.909] 0.334 ± 0.119 [0.276, 0.392]
DFN 0.873 ± 0.043 [0.853, 0.893] 0.926 ± 0.047 [0.904, 0.948] 0.741 ± 0.092 [0.696, 0.786]
GCN 0.871 ± 0.051 [0.847, 0.895] 0.92 ± 0.049 [0.897, 0.942] 0.735 ± 0.107 [0.685, 0.785]
MOTGNN 0.939 ± 0.031 [0.925, 0.953] 0.969 ± 0.023 [0.958, 0.980] 0.872 ± 0.064 [0.842, 0.902]
LGG XGBoost 0.663 ± 0.041 [0.644, 0.682] 0.718 ± 0.032 [0.703, 0.733] 0.661 ± 0.041 [0.642, 0.680]
RF 0.691 ± 0.051 [0.667, 0.715] 0.764 ± 0.044 [0.743, 0.785] 0.668 ± 0.063 [0.638, 0.698]
DFN 0.691 ± 0.053 [0.666, 0.716] 0.759 ± 0.049 [0.736, 0.782] 0.694 ± 0.052 [0.670, 0.718]
GCN 0.692 ± 0.051 [0.668, 0.716] 0.763 ± 0.049 [0.74, 0.785] 0.693 ± 0.057 [0.666, 0.719]
MOTGNN 0.711 ± 0.038 [0.693, 0.729] 0.777 ± 0.046 [0.756, 0.798] 0.710 ± 0.039 [0.692, 0.728]
STAD XGBoost 0.597 ± 0.061 [0.569, 0.625] 0.635 ± 0.070 [0.603, 0.667] 0.542 ± 0.067 [0.511, 0.573]
RF 0.628 ± 0.040 [0.609, 0.647] 0.681 ± 0.052 [0.657, 0.705] 0.559 ± 0.053 [0.534, 0.584]
DFN 0.635 ± 0.035 [0.619, 0.651] 0.691 ± 0.049 [0.668, 0.714] 0.595 ± 0.046 [0.573, 0.617]
GCN 0.64 ± 0.047 [0.618, 0.662] 0.694 ± 0.045 [0.673, 0.715] 0.603 ± 0.053 [0.578, 0.628]
MOTGNN 0.664 ± 0.060 [0.636, 0.692] 0.712 ± 0.069 [0.680, 0.744] 0.627 ± 0.063 [0.597, 0.657]
Note: Results are reported as mean ± standard deviation [95% confidence interval] over 20 runs (n=20). Confidence intervals are calculated using the t-distribution: $\text{CI}=\bar{x}\pm t_{0.975,19}\times\frac{SD}{\sqrt{n}}$ where $n=20$, $t_{0.975,19}=2.093$. All intervals are symmetric about the mean.
Figure 5. Classification performance comparison for the COADREAD, LGG, and STAD datasets. Bars show mean scores; error bars represent 95% confidence intervals over 20 independent runs.
[…]
[Classification Plot]Classification Bar Plots with Error Bars
Figure 3 presents the feature dimensions before and after applying XGBoost within the MOTGNN framework. The reduced omics matrices $X_{1}^{*}$, $X_{2}^{*}$, and $X_{3}^{*}$ represent the selected feature subsets for DNA methylation, mRNA, and miRNA, respectively. Note that the reduced feature sizes ($p_{i}^{*}$) vary across omics types due to the nature of XGBoost’s feature selection: the final selected set is the union of features used across all $M$ decision trees (see details in Section 3.2). Table 4 highlights the structural properties across MOTGNN’s constructed feature graphs ($G_{1}$-$G_{3}$). Figure 4 presents the relative graph importance for the three constructed graphs ($G_{1}$-$G_{3}$). The edge-to-node ratios (2.1-2.8) suggest sparse graph structures, with miRNA ($G_{3}$) exhibiting the highest edge density (2.35-2.79) but the lowest relative importance (24-29%), indicating limited discriminative value. In COADREAD, mRNA ($G_{2}$) dominates (42% importance), while methylation ($G_{1}$) and mRNA ($G_{2}$) contribute equally in LGG and STAD (35-37%). The consistently high methylation importance (34-37%) across datasets underscores its robustness as a biomarker, whereas mRNA’s variability reflects tissue-specific expression dynamics. These results demonstrate MOTGNN’s ability to prioritize informative omics modalities while maintaining computational efficiency through sparse graph structures, with edge counts remaining tractable for clinical-scale applications.
Table 5 presents classification results averaged over 20 independent train-validation-test splits. Our proposed MOTGNN model consistently outperforms all baselines (XGBoost, RF, DFN, and GCN) across all metrics (Accuracy, ROC-AUC, and F1-score) on all three datasets (COADREAD, LGG, and STAD). In COADREAD, MOTGNN achieves an accuracy of 93.9%, outperforming XGBoost (91.0%), RF (80.1%), DFN (87.3%), and GCN (87.1%). The performance gap is even more evident in F1-score: MOTGNN attains 87.2%, marking a 9.3% absolute improvement over XGBoost (77.9%) and a substantial 53.8% advantage over RF (33.4%). Given COADREAD’s imbalanced class distribution (254:78), this improvement highlights MOTGNN’s robustness in handling skewed datasets, where traditional models like RF show substantially reduced performance. The ROC-AUC results further validate MOTGNN’s stability, with scores exceeding 96.9% for COADREAD and maintaining at least 3% improvements over the best baseline across all datasets. While XGBoost performs competitively in COADREAD (96.4%), its inconsistency becomes apparent when examining results across datasets, dropping to 71.8% for LGG and 63.5% for STAD. DFN shows stronger baseline performance, particularly in STAD (69.1% ROC-AUC), but still falls short of MOTGNN’s 71.2%. GCN exhibits moderate performance across datasets and outperforms DFN slightly in LGG and STAD, which suggests some benefit from modeling feature relationships through a graph structure. However, GCN constructs only one graph over concatenated multi-omics features, which may weaken omics-specific signals and introduce noise from irrelevant cross-omics interactions. In contrast, MOTGNN’s modality-specific graph-based architecture provides additional performance gains, particularly under class imbalance, highlighting its suitability for complex biomedical classification tasks.
Figure 6. F1-score comparison across 20 runs on the imbalanced COADREAD dataset. (A) Box plots show median (labeled), interquartile range, and outliers. (B) Violin plots show median (labeled) and illustrate score distributions. MOTGNN demonstrates both high performance and low variability.
[…]
[BoxandViolinPlot]Box and violin plots showing F1-score distributions for different models on COADREAD.
Figure 5 visually compares model performance across all datasets and evaluation metrics, highlighting MOTGNN’s consistent advantage. Although some confidence intervals slightly overlap, MOTGNN consistently achieves higher mean scores. The most significant gaps appear in F1-scores for COADREAD, where MOTGNN (87.2%) far exceeds RF (33.4%), with non-overlapping confidence intervals. The figure also illustrates that, despite XGBoost’s strong performance in COADREAD (91% accuracy), its accuracy drops considerably in LGG and STAD, underscoring its lack of robustness across datasets. Figure 6 highlights model robustness on the imbalanced COADREAD dataset (class ratio 254:78) by comparing F1-scores over 20 independent experiments. MOTGNN achieves the highest median F1-score (0.871), outperforming XGBoost (0.794), DFN (0.771), GCN (0.735), and RF (0.349). The violin plot reveals a tight distribution around MOTGNN’s median, indicating low variability. In contrast, XGBoost, DFN, and GCN show wider interquartile ranges, while RF exhibits low performance and high variance. These results underscore MOTGNN’s strength in handling class imbalance, a key challenge in biomedical data analysis.
Table 6 presents the top 30 most important biomarkers identified by MOTGNN for the COADREAD dataset (Colorectal Adenocarcinoma and Rectal Adenocarcinoma), illustrating the feature-level interpretability of our framework. COADREAD encompasses colorectal cancer in a broader context and is commonly analyzed jointly, as colon and rectal adenocarcinomas share overlapping molecular signatures and clinical features. This dataset was selected due to its well-characterized molecular landscape and the strong classification performance achieved by our model. The same feature-ranking procedure can also be applied to the other two cancer types: STAD (Stomach Adenocarcinoma) and LGG (Lower Grade Glioma). Among the top-ranked features, Secreted Frizzled-Related Protein 4 (SFRP4) is a well-documented biomarker associated with colorectal cancer (Huang et al., 2010; Yu et al., 2022; Frank et al., 2010; Liu et al., 2020). SFRP4 is overexpressed in tumor tissues, and its promoter hypermethylation has been observed as an early event in cancer progression. Functionally, it acts as a modulator of Wnt signaling, thereby influencing cell proliferation and tumor growth. While our current study focuses on computational identification through feature importance scores, biological validation remains an important next step. These selected biomarkers, including SFRP4, serve as potential candidates for further investigation and may contribute to discovering novel diagnostic or therapeutic strategies for diseases.
Table 6. Top 30 biomarkers identified by MOTGNN for COADREAD by omics type (methylation, mRNA, and miRNA)
Omics type Biomarkers
meth LRRC37A2, C9orf70, SNORD98, SFRP4, ELAVL2
HOXD9, HOXC6, DENND2C, MIR519B, DDIT4L
mRNA MUC12—10071, KDELR2—11014, H3F3A—3020
RPL21—6144, AKT1—207, HOXB13—10481
ILVBL—10994, MAEA—10296, PRAC—84366
H3F3C—440093
miRNA hsa-let-7f-2, hsa-let-7g, hsa-mir-10b, hsa-mir-1201
hsa-mir-1270, hsa-mir-1274b, hsa-mir-191
hsa-mir-206, hsa-mir-30c-2, hsa-mir-425
Note: The COADREAD dataset refers to Colorectal Adenocarcinoma and Rectal Adenocarcinoma (See data details in Table 2). Biomarkers are ranked and selected from three omics types: DNA methylation (meth), gene expression (mRNA), and microRNA expression (miRNA).
Table 7 reveals the computational trade-offs associated with these performance improvements. MOTGNN requires the longest training times, averaging 1.75 minutes for COADREAD and up to 2.39 minutes for LGG, compared to GCN’s 0.59-0.94 minutes, DFN’s 0.3-0.39 minutes, XGBoost’s 0.16-0.25 minutes, and RF’s remarkably fast 0.05 minutes per run. However, these computational costs remain practical for real-world applications, especially considering the analysis was conducted using only four CPUs per job. The increased runtime is justified by MOTGNN’s superior and stable performance across diverse datasets and metrics.
Table 7. Computation time comparison across models and datasets (in minutes).
Data Model Time (min)
COADREAD XGBoost 0.25
RF 0.045
DFN 0.385
GCN 0.585
MOTGNN 1.75
LGG XGBoost 0.195
RF 0.05
DFN 0.36
GCN 0.94
MOTGNN 2.385
STAD XGBoost 0.165
RF 0.05
DFN 0.3
GCN 0.805
MOTGNN 1.45
[…]
6. Conclusion
We present MOTGNN, a novel graph neural network framework for multi-omics data integration and disease classification. By leveraging XGBoost for omics-specific supervised graph construction and modality-specific graph neural networks for latent representation learning, MOTGNN effectively captures complex relationships among features that are often overlooked by traditional methods. Applied to real-world cancer datasets, MOTGNN consistently outperforms baseline models across evaluation metrics, demonstrating its strength in integrating heterogeneous omics data and improving predictive performance. In particular, MOTGNN maintains strong robustness under class imbalance, a common challenge in biomedical data analysis. In addition to its predictive power, MOTGNN provides interpretability at both the feature and omics levels. Feature-level importance scores enable the identification of top-ranked biomarkers, which may serve as promising candidates for further biological investigation. Meanwhile, omics-level contributions reveal the relative informativeness of each data modality (e.g., methylation, mRNA, or miRNA), offering valuable insight into their role across different disease types. Future extensions of the MOTGNN framework may include incorporating additional omics modalities, additional cancer types and non-cancer diseases, or adapting the architecture for more complex prediction tasks, such as multi-class classification or survival analysis. Further enhancing interpretability will also be a key direction, especially to better support clinical and biological decision-making.
Data and Code Availability
The harmonized datasets used in our study and the code for our model are available upon request or publication.
Acknowledgements.
This work is partially supported by the National Institute of General Medical Sciences of the National Institutes of Health (NIH/NIGMS) under Award P20GM104420. Any findings, comments, and conclusions expressed in this paper do not reflect the views of NIH.

### v2

@@
 Integrating multi-omics data, such as DNA methylation, mRNA expression, and microRNA (miRNA) expression, offers a comprehensive view of the biological mechanisms underlying disease.
-However, the high dimensionality and complex interactions among omics layers present major challenges for predictive modeling.
+However, the high dimensionality of multi-omics data, the heterogeneity across modalities, and the lack of reliable biological interaction networks make meaningful integration challenging.
+In addition, many existing models rely on handcrafted similarity graphs, are vulnerable to class imbalance, and often lack built-in interpretability, limiting their usefulness in biomedical applications.
 We propose Multi-Omics integration with Tree-generated Graph Neural Network (MOTGNN), a novel and interpretable framework for binary disease classification.
-MOTGNN employs eXtreme Gradient Boosting (XGBoost) to perform omics-specific supervised graph construction, followed by modality-specific Graph Neural Networks (GNNs) for hierarchical representation learning, and a deep feedforward network for cross-omics integration.
-On three real-world disease datasets, MOTGNN outperforms state-of-the-art baselines by 5-10% in accuracy, ROC-AUC, and F1-score, and remains robust to severe class imbalance (e.g., 87.2% vs.
-33.4% F1 on imbalanced data).
-The model maintains computational efficiency through sparse graphs (2.1-2.8 edges per node) and provides built-in interpretability, revealing both top-ranked biomarkers and the relative contributions of each omics modality.
-These results highlight MOTGNN’s potential to improve both predictive accuracy and interpretability in multi-omics disease modeling.
+MOTGNN employs eXtreme Gradient Boosting (XGBoost) for omics-specific supervised graph construction, followed by modality-specific Graph Neural Networks (GNNs) for hierarchical representation learning, and a deep feedforward network for cross-omics integration.
+Across three real-world disease datasets, MOTGNN outperforms state-of-the-art baselines by 5-10% in accuracy, ROC-AUC, and F1-score, and remains robust to severe class imbalance.
+The model maintains computational efficiency through the use of sparse graphs and provides built-in interpretability, revealing both top-ranked biomarkers and the relative contributions of each omics modality.
+These results highlight the potential of MOTGNN to improve both predictive accuracy and interpretability in multi-omics disease modeling.
 Graph neural networks, Multi-omics integration, XGBoost, Disease classification, Model interpretability
-conference: ; ;
+conference: arXiv; arXiv; 2025
 1. Introduction
 A comprehensive understanding of complex diseases, such as cancer, cardiovascular disease, and neurodegenerative disorders, is critical for advancing public health and enabling precision medicine.
-These diseases are not driven by isolated molecular events but emerge from intricate interactions among genomic, epigenomic, and transcriptomic factors (Subramanian et al., 2020).
-Recent advances in high-throughput biotechnologies have enabled the simultaneous collection of different layers of molecular regulation, such as gene expression (messenger RNA), epigenetic modifications (DNA methylation), and non-coding RNA activity (e.g., microRNA), from the same set of biological samples (Hood and Rowen, 2013).
+These diseases are not driven by isolated molecular events but instead emerge from intricate interactions among genomic, epigenomic, and transcriptomic factors (Subramanian et al., 2020).
+Recent advances in high-throughput biotechnologies have enabled the simultaneous collection of multiple molecular layers, such as gene expression (mRNA), epigenetic modifications (DNA methylation), and non-coding RNA activity (e.g., microRNA), from the same set of biological samples (Hood and Rowen, 2013).
 These omics datasets capture intricate relationships among biological entities and molecular processes, offering insights into disease mechanisms.
-While individual omics types provide valuable information, different modalities reflect distinct biological processes and contribute complementary information.
-Integrating multiple omics (multi-omics) data is essential to comprehensively understand the complex, multi-layered systems underlying disease processes and their progression (Waqas et al., 2024).
+While individual omics types provide valuable information, each reflects distinct biological processes and contributes complementary information.
+Integrating multiple omics (multi-omics) data is therefore essential for comprehensively understanding the multi-layered systems underlying disease processes and progression (Waqas et al., 2024).
 However, integrating heterogeneous omics data presents unique challenges.
-A fundamental issue is that the number of samples ($n$) is substantially smaller than the number of features ($p$).
-For example, a study may involve only a few hundred participants while measuring tens of thousands of molecular features.
-This high dimensionality complicates learning and increases the risk of overfitting.
+A fundamental issue is that the number of samples ($n$) is typically much smaller than the number of features ($p$).
+For example, a study may involve only a few hundred participants while measuring tens of thousands of molecular variables.
+This high dimensionality complicates model training and increases the risk of overfitting.
 Moreover, not all measured features are biologically relevant; noisy or redundant features can obscure meaningful patterns.
-Although prior biological knowledge, such as gene-gene or protein-protein interactions, has been identified experimentally, our knowledge is constrained by experimental design and equipment constraints.
-Consequently, there is a growing need for computational models that can reliably infer biologically meaningful structures and interactions from multi-omics data (Jiang et al., 2025).
-Traditional machine learning and deep learning models have been extensively applied to address the high dimensionality and complexity of omics data (Ayman et al., 2023; Zaghlool and Attallah, 2022), but most of these approaches assume a Euclidean data structure.
-Classical machine learning methods, such as random forests (Breiman, 2001) and gradient boosting methods, such as XGBoost (Chen and Guestrin, 2016), operate by constructing ensembles of decision trees and are effective for classification tasks, but they typically treat features as independent and fail to capture structural relationships among them.
-Deep learning models (Lecun et al., 2015), such as deep feedforward networks (DFNs) (Goodfellow et al., 2016), convolutional neural networks, and recurrent neural networks, can model nonlinear feature dependencies but are not designed to process non-Euclidean data structures such as graphs.
-These approaches typically require large training sets to avoid overfitting.
-Moreover, their black-box nature often limits interpretability, which is essential in biomedical applications where transparency and reproducibility are critical.
-Furthermore, concatenating features from different omics modalities as input can obscure modality-specific relationships and introduce additional noise.
-Graph-based approaches offer a natural alternative for modeling complex biological systems.
-In omics applications, nodes can represent features (e.g., genes) or samples, and edges encode various relationships, such as similarity, co-expression, or learned interactions.
-Graph Neural Networks (GNNs) (Bronstein et al., 2017; Wu et al., 2020; Zhang et al., 2020; Zhou et al., 2020; Chen et al., 2023) have emerged as powerful tools for learning from graph-structured data by recursively aggregating information from each node’s neighborhood.
-GNNs capture both local and global graph topology and have found increasing use in biomedical applications, including molecular modeling, medical imaging, drug discovery, and disease prediction (Zhang et al., 2021; Johnson et al., 2024; Waqas et al., 2024; Zhong et al., 2025).
-Foundational models such as message passing neural networks (MPNNs) (Gilmer et al., 2017), graph convolutional networks (GCNs) (Kipf and Welling, 2017), graph attention networks (GATs) (Petar et al., 2018), GraphSAGE (Hamilton et al., 2017), and graph isomorphism networks (GINs) (Xu et al., 2019) have significantly advanced the expressiveness and scalability of GNN-based learning.
+Although biological interaction networks such as gene-gene or protein-protein interactions have been identified experimentally, current knowledge is constrained by limitations in experimental design and equipment.
+Consequently, there is a growing need for computational models that can reliably infer biologically meaningful structures and interactions directly from multi-omics data (Jiang et al., 2025).
+Traditional machine learning and deep learning models have been widely applied to address the high dimensionality and complexity of omics data (Zaghlool and Attallah, 2022; Ayman et al., 2023), but most assume a Euclidean data structure.
+Classical methods such as random forests (Breiman, 2001) and gradient boosting algorithms like XGBoost (Chen and Guestrin, 2016) construct ensembles of decision trees and perform well for classification, yet they typically treat features as independent and fail to capture structural relationships among them.
+Deep learning models (Lecun et al., 2015), including deep feedforward networks (DFNs) (Goodfellow et al., 2016), convolutional neural networks, and recurrent neural networks, can model nonlinear feature dependencies but are not designed to process non-Euclidean data structures such as graphs.
+These approaches typically require large training sets to prevent overfitting and are often criticized for their black-box nature, which limits interpretability, an essential requirement in biomedical research where transparency and reproducibility are critical.
+Furthermore, simply concatenating features from different omics modalities into a single input can obscure modality-specific relationships and introduce additional noise.
+Graph-based approaches provide a natural alternative for modeling complex biological systems.
+In omics applications, nodes can represent features (e.g., genes) or samples, and edges encode various relationships such as similarity, co-expression, or learned interactions.
+Graph Neural Networks (GNNs) (Bronstein et al., 2017; Wu et al., 2020; Zhang et al., 2020; Zhou et al., 2020) have emerged as powerful tools for learning from graph-structured data by recursively aggregating information from each node’s neighborhood.
+GNNs capture both local and global graph topological patterns and have shown increasing success in biomedical domains, including molecular modeling, drug discovery, and disease prediction (Zhang et al., 2021; Johnson et al., 2024; Waqas et al., 2024; Zhong et al., 2025).
+Foundational architectures such as message passing neural networks (MPNNs) (Gilmer et al., 2017), graph convolutional networks (GCNs) (Kipf and Welling, 2017), graph attention networks (GATs) (Petar et al., 2018), GraphSAGE (Hamilton et al., 2017), and graph isomorphism networks (GINs) (Xu et al., 2019) have advanced the expressiveness and scalability of GNN-based learning.
+Recent GNN-based methods have aimed to integrate multi-omics data for disease classification, cancer subtype analysis, and survival prediction.
+MOGONET (Wang et al., 2021) introduces a GCN-based framework for integrating methylation, mRNA, and miRNA through a cross-omics discovery tensor and a view correlation discovery network.
+However, it identifies important features through computationally intensive ablation studies, and its graph construction relies on cosine similarity with a manually tuned edge-to-node ratio.
+MODILM (Zhong et al., 2023) adopts a similar structure but replaces GCN with GAT, while CLCLSA (Zhao et al., 2024) addresses incomplete omics data using cross-omics autoencoders, contrastive learning, and attention mechanisms without explicitly modeling graph structures.
+SUPREME (Kesimoglu and Bozdag, 2023) and MOGAT (Tanvir et al., 2024) extend graph-based modeling to cancer subtype prediction and survival analysis by employing patient similarity networks derived from predefined metrics.
+DeepMoIC (Wu et al., 2024) integrates three omics types using autoencoders, similarity network fusion, and a deep GCN with residual connections, while MoGCN (Li et al., 2022) focuses on breast cancer subtyping with graphs constructed via similarity network fusion.
+Although these models achieve promising results within their respective applications, they primarily rely on unsupervised or distance-based graphs and often lack built-in interpretability or robustness under class imbalance.
 Despite recent advances, applying GNNs to multi-omics data remains challenging and relatively underexplored (Zhang et al., 2021; Valous et al., 2024; Paul et al., 2024; Ballard et al., 2024).
-The irregular nature of graph-structured data further complicates this task, as each node may have a variable number of neighbors, and many standard deep learning operations are not directly applicable in this setting (Chen et al., 2023).
-Although several GNN-based frameworks have been proposed for multi-omics integration, many of these works exhibit substantial limitations, which we summarize into three main areas: (1) Inadequate modeling of high-dimensional, heterogeneous omics data: To address high dimensionality (i.e., the “small $n$, large pp” issue), existing methods typically apply simple filtering techniques, such as removing low-variance or zero-valued features, or selecting features based on univariate statistical tests.
-Some approaches flatten multi-omics data by concatenating omics layers before graph construction, which weakens modality-specific signals and ignores biological heterogeneity.
-Others rely heavily on hand-crafted similarity graphs (e.g., Pearson or cosine similarity), which may miss nonlinear, task-specific relationships, ignore the sparse and scale-free nature of biological graphs, and are sensitive to arbitrary thresholding or preprocessing pipelines.
-(2) Vulnerability to class imbalance: While class imbalance is prevalent in biomedical datasets (e.g., rare diseases, early-stage cancer detections, or minority subtypes), it is rarely addressed during model development or evaluation.
-Many existing models optimize for overall accuracy, which can be misleading when predictions favor the majority class.
-In these scenarios, even high accuracy may reflect poor minority class detection.
-As a result, these approaches often lack robustness and generalizability in imbalanced settings, where accurate prediction of minority cases is of greater biological and clinical importance.
-(3) Limited interpretability and biological insight: Interpretability is essential in biomedical research, as understanding which features or omics modalities drive predictions is essential for trust, transparency, and downstream biological discovery.
-However, most existing GNN-based models lack built-in interpretability.
-Instead, they rely on computationally intensive post hoc analyses or ablation studies to infer importance, which are not scalable and may introduce instability.
-In addition, they typically treat all omics types equally, failing to uncover their relative contributions.
+Existing methods exhibit three key limitations: (1) Inadequate modeling of high-dimensional, heterogeneous omics data: To address the “small $n$, large pp” issue, many methods apply simple filtering (e.g., removing low-variance or zero-valued features) or flatten multi-omics data through concatenation, which weakens modality-specific signals and ignores biological heterogeneity.
+Others rely on handcrafted similarity graphs (e.g., Pearson or cosine), which may miss nonlinear, task-specific relationships and are sensitive to arbitrary thresholding.
+(2) Vulnerability to class imbalance: While class imbalance is common in biomedical datasets (e.g., rare diseases, early-stage cancer detections, or minority subtypes), it is rarely addressed during model development or evaluation.
+Many models emphasize overall accuracy, which can obscure poor minority-class performance and limit generalizability.
+(3) Limited interpretability and biological insight: Interpretability is essential in biomedical research for trust, transparency, and downstream discovery.
+However, most GNN-based models lack built-in interpretability and instead rely on computationally intensive post-hoc analyses or ablation studies that may be unstable.
+Furthermore, they typically treat all omics modalities equally, overlooking their relative contributions to prediction.
 To address these limitations, we propose MOTGNN, a novel and interpretable graph-based framework for multi-omics data integration that combines supervised graph construction with hierarchical representation learning.
-Our key contributions are as follows:
-  • Modality-specific modeling with supervised graph learning: We utilize XGBoost to filter informative features for each omics type, and leverage the structure of the trained decision trees to construct sparse, supervised, and modality-specific graphs.
+Our main contributions are summarized as follows: (1) Modality-specific modeling with supervised graph learning: MOTGNN employs XGBoost to identify informative features for each omics type and uses the trained decision-tree structure to construct sparse, supervised, and modality-specific graphs.
 This approach preserves biologically meaningful interactions while reducing noise and redundancy.
-We then apply GNNs to learn latent representations from each graph, which are integrated through a DFN to model cross-omics interactions.
-  • Imbalance-robust architecture: MOTGNN consistently outperforms existing models across all datasets, achieving 10-50% improvement in F1-score on class-imbalanced real-world data.
-Its design mitigates overfitting to the dominant classes and remains effective in detecting minority class signals, addressing a critical limitation in biomedical classification tasks.
-  • Integrated interpretability at feature and omics levels: MOTGNN enables end-to-end interpretability without requiring post hoc analysis.
-It provides (i) feature-level importance scores to identify top-ranked biomarkers within each modality, and (ii) omics-level contribution scores, offering insights into which data types (e.g., methylation, mRNA, or miRNA) are most influential for prediction.
-This layered interpretability potentially supports downstream biological discovery and facilitates clinical relevance.
-Figure 1.
-Overview of the proposed MOTGNN framework, which involves three key modules: (i) XGBoost for graph construction within each omics modality; (ii) GNN for learning modality-specific embeddings by encoding both feature graphs and associated data matrices; and (iii) DFN that integrates the learned embeddings and performs classification.
-[…]
-[MOTGNN architecture]The MOTGNN framework includes XGBoost for graph construction, a GNN module for encoding, and a DFN for integration and classification.
-2. Related Work
-Many advanced deep learning methods have been developed by incorporating novel model components to analyze omics data.
-For example, MOCAT (Yao et al., 2024) leverages an autoencoder, three omics-specific classifiers, a multi-head attention mechanism, and a confidence network (ConfNet) to integrate and learn from multi-omics data (e.g., methylation, mRNA, and miRNA).
-However, it lacks the ability to capture topological information and thus may miss important structural insights.
-Some GNN-based methods have been developed for single-omics data analysis in disease classification or cancer subtype prediction.
-For example, forgenet (Kong and Yu, 2020) constructs a supervised directed graph using a tree-ensemble model, but it uses the original high-dimensional data without dimension reduction, leading to high computational costs.
-Moreover, it only uses a single omics type, which provides limited information compared to multi-omics integration.
-DMGCN (Yang et al., 2021) integrates multiple biological networks via GCNs, but also focuses on a single omics modality and does not utilize cross-omics information.
-More recent GNN-based methods aim to integrate multi-omics data for disease classification, cancer subtype analysis, and survival prediction.
-MOGONET (Wang et al., 2021) proposes a GCN-based framework to integrate multi-omics data (methylation, mRNA, and miRNA), using a cross-omics discovery tensor and a view correlation discovery network for final label prediction.
-However, it identifies important features through computationally expensive ablation studies.
-In addition, its graph construction is based on cosine similarity with a manually selected threshold determined by a hyperparameter (edge-to-node ratio), which must be tuned.
-MODILM (Zhong et al., 2023) shares a similar architecture with MOGONET but uses GAT instead of GCN.
-CLCLSA (Zhao et al., 2024) handles incomplete multi-omics classification using cross-omics autoencoders to recover missing omics representations, contrastive learning to align representations, and self-attention at both the feature and omics levels for feature selection, but does not explicitly leverage graph structures.
-SUPREME (Kesimoglu and Bozdag, 2023) employs GCNs and patient similarity networks, integrating seven types of omics data for cancer subtype prediction and survival analysis.
-However, it also relies on ablation studies to determine the most effective omics combinations.
-MOGAT (Tanvir et al., 2024) focuses on survival analysis with a similar structure to SUPREME, but uses GAT instead of GCN.
-It conducted integration studies with two to eight omics types and found that using three omics achieved the best performance for breast cancer subtype prediction, suggesting that adding more omics data does not always yield better results.
-However, MOGAT does not address model interpretability, as it does not identify important biomarkers.
-DeepMoIC (Wu et al., 2024) integrates three types of omics data to address cancer subtype classification by combining autoencoders, similarity network fusion, and a deep GCN with residual connections and identity mappings.
-However, it constructs patient similarity networks to guide the learning and lacks omics-wise interpretability.
-MoGCN (Li et al., 2022) targets breast cancer subtype classification using multi-omics data (copy number variation, RPPA, and RNA-seq).
-It applies autoencoders for dimensionality reduction and similarity network fusion for graph construction, and trains a GCN model on the resulting fused network.
-However, its graph construction still relies on distance-based methods.
-While these models demonstrate promising results, several common limitations remain: They require extensive data preprocessing, and the quality of such preprocessing can significantly affect downstream model performance.
-Their graph construction approaches rely on simple similarity metrics, which often fail to capture task-relevant nonlinear interactions and can introduce bias due to arbitrary thresholding.
-They rarely address the class imbalance issue, which is prevalent in biomedical datasets and can significantly impact performance.
-They commonly depend on ablation studies to identify the optimal combination of omics data types, a computationally expensive process that is not easily scalable.
+Omics-specific GNNs learn latent representations from each graph, which are then integrated through a deep feedforward network (DFN) to capture cross-omics interactions.
+(2) Imbalance-robust architecture: MOTGNN consistently outperforms existing models across all datasets, achieving 10-50% improvement in F1-score on class-imbalanced real-world data.
+Its design mitigates overfitting to dominant classes and effectively detects signals from minority classes, addressing a critical limitation in biomedical classification tasks.
+(3) Integrated interpretability at feature and omics levels: MOTGNN provides end-to-end interpretability without requiring post-hoc analysis.
+It generates (i) feature-level importance scores to identify top-ranked biomarkers within each modality, and (ii) omics-level contribution scores to quantify the relative importance of each data type (e.g., methylation, mRNA, or miRNA), supporting both biological discovery and clinical insight.
 2. Methodology
-We summarized the key notations and their associated descriptions in Table 1.
+We summarize the key notations and their associated descriptions in Table 1.
 A graph is denoted as $G(V,E)$, where $V$ represents the set of vertices (or nodes) and $E$ denotes the set of edges.
-Graph structure is encoded by its adjacency matrix $A$, in which the $(i,j)$-th element is one if node $i$ is connected to node $j$, and zero otherwise.
+The graph structure is encoded by its adjacency matrix $A$, where the $(i,j)$-th element is one if node $i$ is connected to node $j$, and zero otherwise.
 For undirected graphs, $A$ is symmetric.
 The adjacency matrix augmented with self-loops is denoted as $\tilde{A}$, where all diagonal elements are set to one to include self-connections.
-Table 1. Key notations and descriptions
+Table 1. Key notations and their descriptions used throughout the paper.
 Notation Description
@@
 $I(\cdot)$ Indicator function
+Figure 1.
+Overview of the proposed MOTGNN framework for multi-omics data integration and disease classification.
+The model comprises three key modules: (i) XGBoost for constructing omics-specific supervised feature graphs; (ii) graph neural network (GNN) for learning modality-specific embeddings by encoding each graph and its corresponding data matrix; and (iii) deep feedforward network (DFN) for integrating the learned embeddings and performing final classification.
 2.1. Proposed Model
-We propose a Multi-Omics integration framework with Tree-generated Graph Neural Network (MOTGNN) model for binary disease classification.
-MOTGNN first leverages eXtreme Gradient Boosting (XGBoost) to select informative features in a supervised manner, as the tree-based feature selection captures strong predictive signals.
-It then utilizes GNNs to learn omics-specific representations from graphs constructed using the structure of trained XGBoost trees.
+We propose MOTGNN, a Multi-Omics integration framework with Tree-generated Graph Neural Networks for binary disease classification.
+MOTGNN first leverages eXtreme Gradient Boosting (XGBoost) to select informative features in a supervised manner, as tree-based feature selection captures strong predictive signals.
+It then utilizes GNNs to learn omics-specific representations from graphs constructed using the structure of the trained XGBoost trees.
 This combination allows the model to extract both local and global patterns from each omics layer while maintaining biological interpretability.
-The framework of MOTGNN is shown in Figure 1, which consists of three primary modules: (1) An XGBoost module for feature graph construction; (2) A GNN module to encode each omics-specific graph and its associated data matrix; and (3) A DFN module to integrate and classify the learned embeddings.
-This modular architecture allows MOTGNN to handle high-dimensional, heterogeneous multi-omics data efficiently while providing interpretable insights at both the feature and omics-layer levels.
-Algorithm 1 demonstrates the steps of the overall structure of MOTGNN.
+The overall framework of MOTGNN (Figure 1) consists of three primary modules: (1) an XGBoost module for feature graph construction; (2) a GNN module to encode each omics-specific graph and its associated data matrix; and (3) a DFN module to integrate and classify the learned embeddings.
+This modular architecture allows MOTGNN to handle high-dimensional, heterogeneous multi-omics data efficiently, while providing interpretable insights at both the feature and omics levels.
+Algorithm 1 outlines the overall procedure of MOTGNN.
 Algorithm 1 MOTGNN
@@
 8:end for
-9:Concatenate $Z_{1}$, $Z_{2}$, and $Z_{3}$ to form a unified representation
-10:Pass the concatenated representation $Z$ into a DFN to perform final binary classification
+9:Concatenate $Z_{1}$, $Z_{2}$, and $Z_{3}$ to form a unified representation $Z=[Z_{1}|Z_{2}|Z_{3}]$
+10:Pass $Z$ into a DFN to perform final binary classification
 11:return Predicted classification labels $\hat{Y}$
@@
 XGBoost (Chen and Guestrin, 2016) is a powerful ensemble method based on gradient-boosted decision trees.
-Compared to traditional gradient boosting machines (GBMs) (Friedman, 2002), XGBoost introduces a regularized learning objective to reduce overfitting and improve generalization.
-In MOTGNN, we employ three XGBoost models, each trained independently to perform binary classification on the omics-specific datasets $X_{1}$, $X_{2}$, and $X_{3}$ alongside the binary response $Y$.
-Unlike random forest, which generates independent trees, each trained XGBoost model generates $M$ sequential decision trees, where every tree improves upon the residuals of the previous ones.
+Compared with traditional gradient boosting machines, it introduces a regularized learning objective to reduce overfitting and improve generalization.
+In MOTGNN, we employ three XGBoost models, each trained independently to perform binary classification on the omics-specific datasets $X_{1}$, $X_{2}$, and $X_{3}$ together with the binary response $Y$.
+Unlike random forests, which generate independent trees, each XGBoost model generates $M$ sequential decision trees, where every tree improves upon the residuals of the previous ones.
 During training, each tree selects a subset of features for node splits.
 We collect the union of all features used across the $M$ trees and define these as the selected features.
-This yields a supervised, tree-guided feature selection process that reduces the original feature dimensions from $p_{1}$, $p_{2}$, and $p_{3}$ to $p_{1}^{*}$, $p_{2}^{*}$, and $p_{3}^{*}$, respectively, where $p_{i}^{*}<p_{i}$ and $p_{1}^{*}\neq p_{2}^{*}\neq p_{3}^{*}$.
-The reduced dimensions preserve features with higher predictive importance, facilitating downstream graph construction and neural modeling.
-Inspired by forgeNet (Kong and Yu, 2020), which constructs feature graphs based on the structure of ensemble trees, we adapt this approach to build omics-specific undirected graphs from XGBoost models.
-Specifically, for each omics dataset, we treat each trained decision tree as a graph, where features involved in splits are nodes, and parent-child relationships form the undirected edges.
-Taking $X_{1}\in\mathbb{R}^{n\times p_{1}}$ along with the binary outcome $Y$ as an example, suppose the XGBoost model generates $M$ sequential decision trees.
-Each tree is considered an undirected graph $G^{m}=(V^{m},E^{m})$ for $m=1,2,\ldots,M$.
-We then take the union of all tree-level graphs to form the aggregated graph $G_{1}=(V_{1},E_{1})$ as:
+This yields a supervised, tree-guided feature selection process that reduces the original feature dimensions from $p_{1}$, $p_{2}$, and $p_{3}$ to $p_{1}^{*}$, $p_{2}^{*}$, and $p_{3}^{*}$, respectively, where $p_{i}^{*}<p_{i}$ and, in general, $p_{1}^{*}\neq p_{2}^{*}\neq p_{3}^{*}$.
+The reduced feature spaces preserve variables with higher predictive importance, facilitating downstream graph construction and neural modeling.
+Inspired by forgeNet (Kong and Yu, 2020), which constructs feature graphs based on the structure of ensemble trees, we adapt this idea to build omics-specific undirected graphs from XGBoost models.
+For each omics dataset, we treat each trained decision tree as a graph, where features involved in splits are represented as nodes and parent-child relationships form undirected edges.
+For example, consider $X_{1}\in\mathbb{R}^{n\times p_{1}}$ and the binary outcome $Y$: if the XGBoost model generates $M$ sequential decision trees, each tree is viewed as an undirected graph $G^{m}=(V^{m},E^{m})$ for $m=1,2,\ldots,M$.
+We then take the union of all tree-level graphs to form an aggregated graph as
 $G_{1}(V_{1},E_{1})=G_{1}(\bigcup_{m=1}^{M}V^{m},\bigcup_{m=1}^{M}E^{m}),$
-where the number of nodes in $G_{1}$ is $|V_{1}|=p_{1}^{*}$, and $p_{1}^{*}<p_{1}$.
-We also add self-loops to the adjacency matrices to preserve self-information in GNN message passing.
-A similar procedure generates $G_{2}$ and $G_{3}$ for the remaining two omics datasets.
+where the number of nodes is $|V_{1}|=p_{1}^{*}$, and $p_{1}^{*}<p_{1}$.
+Self-loops are added to the adjacency matrices to preserve self-information in GNN message passing.
+The similar procedure is used to generate $G_{2}$ and $G_{3}$ for the remaining two omics datasets.
 2.3. Graph Neural Network and Feature Fusion
-We utilize the graph-embedded deep feedforward network (GEDFN) model (Kong and Yu, 2018) as the core GNN component used in MOTGNN.
-Unlike standard fully connected architectures, GEDFN directly incorporates the adjacency matrix into the weight connections between the input and the first hidden layers.
-This design enables the model to respect the feature-level relationships encoded in the graph while enforcing the sparsity, a common assumption for omics-based feature graphs.
-The number of neurons in the first hidden layer must be the same as the input dimension.
-Formally, given the reduced input data $X^{*}\in\mathbb{R}^{n\times p^{*}}$ and its graph adjacency matrix $\tilde{A}\in\mathbb{R}^{p^{*}\times p^{*}}$ (augmented with self-loops), the first hidden layer $Z_{1}\in\mathbb{R}^{n\times p^{*}}$ is computed as
-$Z_{1}=\sigma(X^{*}(W_{in}\odot\tilde{A})+b_{in}),$
-where $W_{in}\in\mathbb{R}^{p^{*}\times p^{*}}$ and $b_{in}\in\mathbb{R}^{n\times p^{*}}$ are the weight and bias matrices, $\odot$ denotes element-wise (Hadamard) multiplication, and $\sigma(\cdot)$ is the activation function.
+We use the graph-embedded deep feedforward network (GEDFN) model (Kong and Yu, 2018) as the core GNN component used in MOTGNN.
+Unlike standard fully connected architectures, GEDFN directly incorporates the adjacency matrix into the weight connections between the input and the first hidden layer.
+This design allows the model to respect the feature-level relationships encoded in the graph while enforcing sparsity, which is a common assumption for omics-based feature graphs.
+The number of neurons in the first hidden layer is set equal to the input dimension.
+Given a reduced input data $X^{*}\in\mathbb{R}^{n\times p^{*}}$ and its augmented graph adjacency matrix $\tilde{A}\in\mathbb{R}^{p^{*}\times p^{*}}$, the first hidden layer $Z_{1}\in\mathbb{R}^{n\times p^{*}}$ is computed as
+$Z_{1}=\sigma(X^{*}(W_{\text{in}}\odot\tilde{A})+b_{\text{in}}),$
+where $W_{\text{in}}\in\mathbb{R}^{p^{*}\times p^{*}}$ and $b_{\text{in}}\in\mathbb{R}^{n\times p^{*}}$ are the weight and bias matrices, $\odot$ denotes element-wise (Hadamard) multiplication, and $\sigma(\cdot)$ is the activation function.
+This ensures that only graph-supported connections are active between the input features and the first hidden layer.
 The design of GEDFN shares conceptual similarities with the graph attention network (GAT) formulation discussed in (Chen et al., 2023), where the GAT layer is represented as
-$Z_{1}=\sigma((W_{attention}\odot A)X+b_{in}),$
-with $W_{attention}\in\mathbb{R}^{n\times n}$ and the nodes representing samples rather than features.
+$Z_{1}=\sigma((W_{\text{attention}}\odot A)X+b_{\text{in}}),$
+with $W_{\text{attention}}\in\mathbb{R}^{n\times n}$ and the nodes representing samples rather than features.
 In this formulation, GAT dynamically learns attention weights over neighbors.
-Similarly, GEDFN enables dynamic weight learning while ensuring that connections between features and the first hidden layer are strictly supported by the feature graph.
+In GEDFN, the graph structure constrains which connections are allowed, while the model still learns connection strengths.
 This yields a sparse and interpretable architecture tailored to omics data.
-Following the training of three omics-specific GNNs, each on its respective dimension-reduced input and constructed graph, we obtain three last hidden representations: $Z_{1}$, $Z_{2}$, and $Z_{3}$.
+After training three omics-specific GNNs, each on its own dimension-reduced input and constructed graph, we obtain three final hidden representations: $Z_{1}$, $Z_{2}$, and $Z_{3}$.
 These embeddings are concatenated to form a unified representation:
 $Z=[Z_{1}|Z_{2}|Z_{3}],$
-where $Z$ denotes the concatenated neurons, and is then passed to a DFN, which consists of fully connected layers and a final softmax output to perform binary disease classification.
+which is then passed to a DFN consisting of fully connected layers and a final softmax output layer to perform binary disease classification.
 Table 2.
-Data description of three TCGA cancer datasets (COADREAD, LGG, and STAD) regarding sample sizes, class distributions, the number of original features, and the number of preprocessed features for each omics type (methylation, mRNA, and miRNA).
+Summary of three TCGA cancer datasets (COADREAD, LGG, and STAD), including sample sizes, class distributions, the number of original features, and the number of preprocessed features for each omics type: DNA methylation (meth), gene expression (mRNA), and microRNA expression (miRNA).
+The preprocessed data were obtained from the Broad GDAC Firehose.
 Abbr. Data Name Samples Class Balance Original Features Preprocessed Features
@@
 STAD Stomach Adenocarcinoma 371 ADC:IAC = 205:166 20,101:20,531:507 2,000:2,000:507
-Note: meth = DNA methylation, mRNA = gene expression, miRNA = microRNA expression. The processed data were obtained from the Broad GDAC Firehose (Lu et al., 2023).
 Figure 2.
-The histogram of methylation, mRNA, and miRNA on preprocessed COADREAD, LGG, and STAD datasets.
-Each omics dataset was individually scaled to [0, 1].
-Different omics types demonstrate distinct distribution patterns.
-[…]
-[Omics distribution]The histograms of methylation, mRNA, and miRNA on COADREAD, LGG, and STAD datasets.
-Each omics dataset was individually scaled to the [0, 1] range using min-max normalization.
-Different omics data demonstrate distinct distribution patterns.
+Distribution of DNA methylation, mRNA, and miRNA features across the preprocessed COADREAD, LGG, and STAD datasets.
+Each omics dataset was independently scaled to the range of [0, 1] using min-max normalization.
+The distinct distributional patterns highlight the heterogeneous characteristics of different omics types.
 2.4. Interpretability on Features and Graphs
-To interpret model predictions and evaluate the contribution of features and graphs, we adopt and extend the connection weights method (Olden and Jackson, 2002).
-Feature importance is assessed by summing the absolute weights connecting each input neuron to the first hidden layer, while accounting for adjacency-based sparsity.
-For feature $j$ in $X_{i}$, the importance score $IF_{j}^{i}$ is defined as:
-$IF_{j}^{i}=\sum_{u=1}^{p_{i}}\left|W_{ju}^{(in)}I(\tilde{A}_{ju}^{i}=1)\right|,$
-where $W^{(\text{in})}$ is the input weight matrix and $I(\cdot)$ is the indicator function.
-After training the model, each input feature is assigned an importance score.
-These scores can be used to rank the features, allowing the selection of top biomarkers for further biological investigation.
-To evaluate the relative importance of the graph for each omics modality, we compute the $L_{1}$ norm of the weight matrix connecting the GNN’s last hidden representations ($Z_{i}$) to the DFN layer.
-The relative graph importance (RIG) is then:
+To interpret model predictions and evaluate the contribution of both features and graphs, we adopt and extend the connection weights method (Olden and Jackson, 2002).
+Feature importance is quantified by summing the absolute values of weights connecting each input neuron to the first hidden layer, while accounting for adjacency-based sparsity in the feature graph.
+For feature $j$ in omics modality $X_{i}$, the importance score $IF_{j}^{i}$ is defined as:
+$IF_{j}^{i}=\sum_{u=1}^{p_{i}}\left|W_{ju}^{(\text{in})}I(\tilde{A}_{ju}^{i}=1)\right|,$
+where $i=1,2,3$, $W^{(\text{in})}$ is the input weight matrix and $I(\cdot)$ is the indicator function.
+The final importance score for feature $j$ across all omics modalities is obtained by summing its importance values:
+$IF_{j}=IF_{j}^{1}+IF_{j}^{2}+IF_{j}^{3}.$
+After training the model, each input feature is assigned an importance score that can be used to rank and identify top biomarkers for downstream biological investigation.
+To assess the relative importance of each omics-specific graph, we compute the $L_{1}$ norm of the weight matrix connecting the GNN’s last hidden representations ($Z_{i}$) to the DFN layer.
+The relative graph importance ($RIG_{i}$) is defined as:
 $RIG_{i}=\frac{\|W_{Z_{i}\leftrightarrow f}\|_{1}}{\sum_{i=1}^{3}\|W_{Z_{i}\leftrightarrow f}\|_{1}},$
-where $i=1,2,3$, $W_{Z_{i}\leftrightarrow f}$ is the weight matrix connecting $Z_{i}$ and DFN, and $\|\cdot\|_{1}$ is the $L_{1}$ norm that sums the absolute values of its inside matrix.
-The total contribution sums to one across all three graphs.
+where $W_{Z_{i}\leftrightarrow f}$ denotes the weight matrix connecting $Z_{i}$ to the DFN, and $\|\cdot\|_{1}$ is the $L_{1}$ norm (the sum of absolute values).
+The total relative importance across all three graphs sums to one.
 2.5. Technical Details
-We implement the xgboost Python library (Chen and Guestrin, 2016) for the XGBoost module in MOTGNN.
-We tuned the number of estimators (determine trees).
-We employ the Tensorflow library (Abadi et al., 2016) for training GNN and DFN modules.
+We implement the XGBoost module in MOTGNN using the xgboost Python library and tune the number of estimators (i.e., the number of trees).
+The GNN and DFN modules are implemented in the Tensorflow library (Abadi et al., 2016).
 Model optimization uses the Adam optimizer (Kingma and Ba, 2014) with mini-batch training.
@@
 For classification, the final output layer uses a softmax function, and the loss is computed using binary cross-entropy:
-$L(Y,\hat{Y})=-\frac{1}{n}\sum_{i=1}^{n}\left\{y_{i}\ln(\hat{y}_{i})+(1-y_{i})\ln(1-\hat{y}_{i})\right\}.$
-We tune hyperparameters across three main categories: (1) architectural (number of layers and hidden neurons), (2) training-related (learning rate, batch size, and epochs), and (3) regularization-related (dropout and L2 penalty).
-To avoid overfitting, we apply strategies including dropout, batch normalization, L2 regularization, and early stopping.
-Data are split into training, validation, and test sets in a 60%:20%:20% ratio, with stratified sampling to address class imbalance.
-We have conducted 20 independent and reproducible train-validation-test splits.
-Hyperparameters are tuned via grid search in a reasonable space and based on experience.
+$L(Y,\hat{Y})=-\frac{1}{n}\sum_{i=1}^{n}\left[y_{i}\ln(\hat{y}_{i})+(1-y_{i})\ln(1-\hat{y}_{i})\right],$
+where $y_{i}\in\{0,1\}$ is the ground truth label, and $\hat{y}_{i}$ is the predicted probability of the positive class.
+We tune hyperparameters across three main categories: (1) architectural (number of layers and hidden neurons), (2) training-related (learning rate, batch size, and epochs), and (3) regularization-related (dropout and $L_{2}$ penalty).
+The $L_{2}$ regularization term is expressed as $L_{\text{reg}}=\lambda\sum_{j}w_{j}^{2}$, where $\lambda$ is the regularization coefficient and $w_{j}$ denotes model weights.
+To avoid overfitting, we apply various strategies including dropout, batch normalization, $L_{2}$ regularization, and early stopping.
+Data are split into training, validation, and test sets in a 60%:20%:20% ratio, with stratified sampling used to address class imbalance.
+We perform 20 independent and reproducible train-validation-test splits.
+Hyperparameters are tuned via a grid search within a reasonable space and informed by prior experience.
 3. Real Data Applications
@@
 The processed multi-omics data and label annotations were obtained from the Broad GDAC Firehose (https://gdac.broadinstitute.org/).
-Each dataset includes matched DNA methylation ($X_{1}$), mRNA expression ($X_{2}$), and miRNA expression ($X_{3}$) data.
-Only samples with complete data across all three omics types were included in our analysis.
+Each dataset includes matched DNA methylation ($X_{1}$), mRNA expression ($X_{2}$), and miRNA expression ($X_{3}$), along with class labels ($Y$).
+Only samples with complete measurements across all three omics types were included in our analysis.
 The COADREAD dataset combines COlorectal ADenocarcinoma (COAD) and REctal ADenocarcinoma (READ) into a binary classification task, with 254 COAD and 78 READ samples.
 The LGG dataset involves grade classification for low-grade gliomas, with 255 Grade 2 and 269 Grade 3 samples.
-The STAD dataset targets the classification of STomach ADenocarcinoma subtypes, with 205 samples from adenocarcinoma (ADC) and 166 from intestinal adenocarcinoma (IAC).
-Among the three datasets, COADREAD has the most imbalanced class distribution.
-Table 2 summarizes the datasets regarding sample sizes, class distributions, the number of original features, and the number of preprocessed features for each omics type.
-Data preprocessing includes quality control steps to remove noisy and redundant features, such as filtering out zero-expression signals, low-variance features, and handling multiple comparisons (Lu et al., 2023).
+The STAD dataset targets the classification of STomach ADenocarcinoma subtypes, with 205 adenocarcinoma (ADC) and 166 intestinal adenocarcinoma (IAC) samples.
+Among the three datasets, COADREAD exhibits the most imbalanced class distribution.
+Table 2 summarizes the datasets in terms of sample sizes, class distributions, and the number of original and preprocessed features for each omics type.
+Data preprocessing included standard quality control steps to remove noisy and redundant features, such as filtering out zero-expression signals and low-variance features, as well as adjusting for multiple comparisons (Lu et al., 2023).
 Each omics dataset was independently normalized to the range [0, 1] using min-max scaling.
-The preprocessed multi-omics data shared in (Lu et al., 2023) contains minor inconsistencies in sample sizes across omics modalities.
-For the COADREAD dataset, meth and mRNA data include 332 samples, while miRNA contains 337 samples.
-We addressed this discrepancy by removing the five extra miRNA samples to ensure alignment.
-Similarly, in the STAD dataset, meth and mRNA data comprise 371 samples, whereas miRNA includes 372 samples.
-Here, we excluded one redundant miRNA sample to maintain consistency.
-The histograms in Figure 2 show the distribution of feature values across three omics types (methylation, mRNA, and miRNA) for the COADREAD, LGG, and STAD datasets.
-The distinct distributional patterns observed across omics types within each dataset highlight each modality’s inherent heterogeneity and diverse data characteristics.
-These differences across the three datasets help evaluate the potential of MOTGNN’s generalizability.
+The preprocessed data shared in (Lu et al., 2023) contained minor inconsistencies in sample sizes across omics types.
+For COADREAD, methylation and mRNA data each included 332 samples, while miRNA had 337.
+We removed the five extra miRNA samples to ensure alignment.
+Similarly, for STAD, methylation and mRNA data each had 371 samples, whereas miRNA had 372; we excluded one redundant miRNA sample for consistency.
+The histograms in Figure 2 show the feature distributions across three omics types (methylation, mRNA, and miRNA) for COADREAD, LGG, and STAD.
+The distinct patterns among omics types highlight the inherent heterogeneity and complementary nature of each modality.
+These differences across datasets help evaluate the generalizability of MOTGNN.
 3.2. Comparison and Evaluation Metrics
-We evaluate the performance of our proposed model, MOTGNN, in comparison to several baseline machine learning and deep learning models: XGBoost (XGB) (Chen and Guestrin, 2016), random forest (RF) (Breiman, 2001), deep feedforward network (DFN) (Goodfellow et al., 2016), and graph convolutional network (GCN) (Kipf and Welling, 2017).
-The number of trees for XGBoost and RF is both at the default, which is 100.
-The number of hidden layers for the DFN and GCN models is two, with several hidden neurons of 64 per layer.
-All baseline models input the column-wise concatenation of the three preprocessed omics data types (methylation, mRNA, and miRNA).
-Table 3 summarizes the tuning ranges and selected values (optimal choices in bold).
+We evaluate the performance of our proposed model, MOTGNN, which by default employs XGBoost to generate three graphs (one per omics type) and uses GEDFN as the main graph neural network (GNN) component.
+The model is compared with four baseline machine learning and deep learning models: XGBoost, random forest (RF), deep feedforward network (DFN), and graph convolutional network (GCN).
+To assess the contribution of individual architectural components, we also design two ablation variants of MOTGNN: (1) $\mathrm{MOTGNN}_{\text{gcn}}$ replaces the GEDFN module with a standard GCN, and (2) $\mathrm{MOTGNN}_{\text{rf}}$ uses RF instead of XGBoost for graph generation.
+These ablation models share the same overall architecture as MOTGNN but differ only in the modified component.
+The number of trees for both XGBoost and RF is set to 100.
+The DFN and GCN models each have two hidden layers with 64 neurons per layer.
+For XGBoost, RF, DFN, and GCN, the input consists of the column-wise concatenation of the three preprocessed omics data types (methylation, mRNA, and miRNA).
+In contrast, $\mathrm{MOTGNN}_{\text{gcn}}$ and $\mathrm{MOTGNN}_{\text{rf}}$ process each omics modality separately, using three parallel GNNs, one for each omics type.
+The graphs used in GCN and $\mathrm{MOTGNN}_{\text{gcn}}$ are generated by XGBoost, while $\mathrm{MOTGNN}_{\text{rf}}$ are generated by RF.
+Table 3 summarizes the hyperparameter tuning ranges and selected values (optimal choices in bold) of MOTGNN.
 We report three standard evaluation metrics: accuracy, ROC-AUC, and F1-score.
 Accuracy measures the proportion of correctly classified samples.
-ROC-AUC (area under the receiver operating characteristic curve) (Green et al., 1966) quantifies the model’s ability to distinguish between classes.
-The F1-score is defined as $F1=\frac{2\times\text{precision}\times\text{recall}}{\text{precision}+\text{recall}},$ where precision refers to the positive predictive value, and recall corresponds to sensitivity or true positive rate.
-As the harmonic mean of precision and recall, the F1-score provides a balanced evaluation, especially in scenarios with imbalanced class distributions, and mitigates bias from simply predicting the majority class.
-These metrics comprehensively assess classification performance across different datasets and modeling approaches.
-We conducted all computations on the Falcon supercomputer (Idaho C3+3 Collaboration, 2022), requesting four CPUs and 20 GB of memory per job.
-Table 3. Hyperparameter tuning for XGBoost and neural networks components in MOTGNN on COADREAD, LGG, and STAD datasets. The optimal choices are in bold.
+ROC-AUC (area under the receiver operating characteristic curve) quantifies the model’s ability to distinguish between classes.
+F1-score is defined as
+$F1=\frac{2\times\text{precision}\times\text{recall}}{\text{precision}+\text{recall}},$
+where precision refers to the positive predictive value and recall corresponds to sensitivity (true positive rate).
+As the harmonic mean of precision and recall, the F1-score provides a balanced assessment under imbalanced class distributions, mitigating bias toward the majority class.
+All computations were conducted on the Falcon supercomputer (Idaho C3+3 Collaboration, 2022), using four CPUs and 20 GB of memory per job.
+Table 3. Hyperparameter tuning ranges for the XGBoost and neural network components in MOTGNN across the COADREAD, LGG, and STAD datasets. Optimal values are highlighted in bold.
 Hyperparameter Tuning Range
@@
 Dropout Rate 0.2, 0.3, 0.5
-L2 Regularization Lambda 0.01
+$L_{2}$ Regularization Coefficient ($\lambda$) 0.01
 Early Stopping Patience 5, 10
@@
 Figure 3.
-Feature dimensions before and after XGBoost on COADREAD, LGG, and STAD datasets.
-The bar plots compare the preprocessed feature dimensions (pre-selection) against the reduced dimensions (post-selection).
-[…]
-[Dimension reduction]Feature Dimension before and after XGBoost selection
-5. Results and Discussion
-Table 4. Structural properties of sparse graphs constructed by MOTGNN for COADREAD, LGG, and STAD datasets.
+Feature dimensions before and after XGBoost-based selection on COADREAD, LGG, and STAD datasets.
+The bar plots compare preprocessed feature dimensions (pre-selection) with reduced dimensions (post-selection), showing substantial dimensionality reduction across all omics types.
+Table 4.
+Structural properties of sparse graphs constructed by MOTGNN for COADREAD, LGG, and STAD datasets.
+$E_{i}$ denotes the number of edges (including self-loops) in graph $G_{i}$, and $m_{i}$ represents the edge-to-node ratio, $m_{i}=E_{i}/p_{i}^{*}$.
 Dataset Edges Edge/Node
@@
 STAD 674:646:639 2.17:2.10:2.61
-Note: The number of edges includes the self-loops, $E_{i}$ is the set of edges for Graph $G_{i}$, $m_{i}$ is the edge to node ratio with $m_{i}=\frac{E_{i}}{p_{i}^{*}}$.
-Figure 4.
-Relative graph importance of different omics types (methylation, mRNA, and miRNA) in MOTGNN’s classification performance across COADREAD, LGG, and STAD datasets.
-Bar heights represent each omics layer’s contribution, revealing dataset-specific biological dominance.
-[…]
-[Relative graph importance]Relative Graph Importance of different omics type (meth, mRNA, and miRNA) for Classification Task on COADREAD, LGG, and STAD datasets
-Note: $G_{i}$ = Graphs from reduced features (methylation:mRNA:miRNA), $RIG_{i}$ = Relative importance of graph $G_{i}$ in MOTGNN (normalized to sum to 1.0 per dataset).
 Table 5.
-Classification performance comparison on COADREAD, LGG, and STAD datasets across different evaluation metrics (Mean ± SD [95% CI]).
-MOTGNN consistently outperforms baseline models across all datasets, highlighting its robustness in integrating multi-omics data (Best results are bolded)
+Classification performance comparison of MOTGNN and other models on COADREAD, LGG, and STAD datasets.
+Results are reported as mean ± standard deviation [95% confidence interval] over 20 independent runs.
+Baseline models include XGBoost, RF (random forest), DFN (deep feedforward network), and GCN (graph convolutional network).
+$\mathrm{MOTGNN}_{\text{gcn}}$ and $\mathrm{MOTGNN}_{\text{rf}}$ represent MOTGNN variants that respectively replace GEDFN with GCN or use RF instead of XGBoost for graph generation.
+Bold values indicate the best performance.
 Dataset Model Accuracy ROC-AUC F1-Score
@@
 DFN 0.873 ± 0.043 [0.853, 0.893] 0.926 ± 0.047 [0.904, 0.948] 0.741 ± 0.092 [0.696, 0.786]
-GCN 0.871 ± 0.051 [0.847, 0.895] 0.92 ± 0.049 [0.897, 0.942] 0.735 ± 0.107 [0.685, 0.785]
+GCN 0.871 ± 0.051 [0.847, 0.895] 0.920 ± 0.049 [0.897, 0.942] 0.735 ± 0.107 [0.685, 0.785]
+$\mathrm{MOTGNN}_{\text{gcn}}$ 0.895 ± 0.039 [0.877, 0.913] 0.928 ± 0.034 [0.912, 0.944] 0.783 ± 0.078 [0.746, 0.820]
+$\mathrm{MOTGNN}_{\text{rf}}$ 0.933 ± 0.028 [0.920, 0.946] 0.967 ± 0.026 [0.955, 0.979] 0.859 ± 0.061 [0.830, 0.888]
 MOTGNN 0.939 ± 0.031 [0.925, 0.953] 0.969 ± 0.023 [0.958, 0.980] 0.872 ± 0.064 [0.842, 0.902]
@@
 DFN 0.691 ± 0.053 [0.666, 0.716] 0.759 ± 0.049 [0.736, 0.782] 0.694 ± 0.052 [0.670, 0.718]
-GCN 0.692 ± 0.051 [0.668, 0.716] 0.763 ± 0.049 [0.74, 0.785] 0.693 ± 0.057 [0.666, 0.719]
+GCN 0.692 ± 0.051 [0.668, 0.716] 0.763 ± 0.049 [0.740, 0.785] 0.693 ± 0.057 [0.666, 0.719]
+$\mathrm{MOTGNN}_{\text{gcn}}$ 0.696 ± 0.050 [0.673, 0.719] 0.764 ± 0.047 [0.742, 0.786] 0.692 ± 0.052 [0.668, 0.716]
+$\mathrm{MOTGNN}_{\text{rf}}$ 0.706 ± 0.030 [0.692, 0.720] 0.777 ± 0.038 [0.759, 0.795] 0.703 ± 0.036 [0.686, 0.720]
 MOTGNN 0.711 ± 0.038 [0.693, 0.729] 0.777 ± 0.046 [0.756, 0.798] 0.710 ± 0.039 [0.692, 0.728]
@@
 DFN 0.635 ± 0.035 [0.619, 0.651] 0.691 ± 0.049 [0.668, 0.714] 0.595 ± 0.046 [0.573, 0.617]
-GCN 0.64 ± 0.047 [0.618, 0.662] 0.694 ± 0.045 [0.673, 0.715] 0.603 ± 0.053 [0.578, 0.628]
+GCN 0.640 ± 0.047 [0.618, 0.662] 0.694 ± 0.045 [0.673, 0.715] 0.603 ± 0.053 [0.578, 0.628]
+$\mathrm{MOTGNN}_{\text{gcn}}$ 0.653 ± 0.062 [0.624, 0.682] 0.686 ± 0.057 [0.659, 0.713] 0.613 ± 0.076 [0.577, 0.649]
+$\mathrm{MOTGNN}_{\text{rf}}$ 0.657 ± 0.048 [0.635, 0.679] 0.710 ± 0.049 [0.687, 0.733] 0.616 ± 0.058 [0.589, 0.643]
 MOTGNN 0.664 ± 0.060 [0.636, 0.692] 0.712 ± 0.069 [0.680, 0.744] 0.627 ± 0.063 [0.597, 0.657]
-Note: Results are reported as mean ± standard deviation [95% confidence interval] over 20 runs (n=20).
-Confidence intervals are calculated using the t-distribution: $\text{CI}=\bar{x}\pm t_{0.975,19}\times\frac{SD}{\sqrt{n}}$ where $n=20$, $t_{0.975,19}=2.093$.
-All intervals are symmetric about the mean.
-Figure 5. Classification performance comparison for the COADREAD, LGG, and STAD datasets. Bars show mean scores; error bars represent 95% confidence intervals over 20 independent runs.
-[…]
-[Classification Plot]Classification Bar Plots with Error Bars
+Figure 4.
+Comparison of classification performance across COADREAD, LGG, and STAD datasets.
+Bars show mean scores, and error bars represent 95% confidence intervals over 20 independent runs.
+MOTGNN consistently achieves the highest scores across accuracy, ROC-AUC, and F1 metrics.
+Figure 5.
+F1-score comparison across 20 independent runs on the imbalanced COADREAD dataset (class ratio 254:78).
+(A) Box plots display the median (labeled), interquartile range, and outliers.
+(B) Violin plots illustrate the distribution of F1-scores with labeled medians.
+MOTGNN achieves the highest median score with the lowest variability, demonstrating stable and robust performance under class imbalance.
+Figure 6.
+Relative importance of feature graphs from different omics types (methylation, mRNA, and miRNA) in MOTGNN across COADREAD, LGG, and STAD datasets.
+Bar heights indicate the contribution of each omics layer to the overall classification performance, revealing dataset-specific biological dominance patterns.
+4. Results and Discussion
+Feature Selection and Graph Properties.
 Figure 3 presents the feature dimensions before and after applying XGBoost within the MOTGNN framework.
 The reduced omics matrices $X_{1}^{*}$, $X_{2}^{*}$, and $X_{3}^{*}$ represent the selected feature subsets for DNA methylation, mRNA, and miRNA, respectively.
-Note that the reduced feature sizes ($p_{i}^{*}$) vary across omics types due to the nature of XGBoost’s feature selection: the final selected set is the union of features used across all $M$ decision trees (see details in Section 3.2).
-Table 4 highlights the structural properties across MOTGNN’s constructed feature graphs ($G_{1}$-$G_{3}$).
-Figure 4 presents the relative graph importance for the three constructed graphs ($G_{1}$-$G_{3}$).
-The edge-to-node ratios (2.1-2.8) suggest sparse graph structures, with miRNA ($G_{3}$) exhibiting the highest edge density (2.35-2.79) but the lowest relative importance (24-29%), indicating limited discriminative value.
-In COADREAD, mRNA ($G_{2}$) dominates (42% importance), while methylation ($G_{1}$) and mRNA ($G_{2}$) contribute equally in LGG and STAD (35-37%).
-The consistently high methylation importance (34-37%) across datasets underscores its robustness as a biomarker, whereas mRNA’s variability reflects tissue-specific expression dynamics.
-These results demonstrate MOTGNN’s ability to prioritize informative omics modalities while maintaining computational efficiency through sparse graph structures, with edge counts remaining tractable for clinical-scale applications.
-Table 5 presents classification results averaged over 20 independent train-validation-test splits.
-Our proposed MOTGNN model consistently outperforms all baselines (XGBoost, RF, DFN, and GCN) across all metrics (Accuracy, ROC-AUC, and F1-score) on all three datasets (COADREAD, LGG, and STAD).
-In COADREAD, MOTGNN achieves an accuracy of 93.9%, outperforming XGBoost (91.0%), RF (80.1%), DFN (87.3%), and GCN (87.1%).
-The performance gap is even more evident in F1-score: MOTGNN attains 87.2%, marking a 9.3% absolute improvement over XGBoost (77.9%) and a substantial 53.8% advantage over RF (33.4%).
-Given COADREAD’s imbalanced class distribution (254:78), this improvement highlights MOTGNN’s robustness in handling skewed datasets, where traditional models like RF show substantially reduced performance.
-The ROC-AUC results further validate MOTGNN’s stability, with scores exceeding 96.9% for COADREAD and maintaining at least 3% improvements over the best baseline across all datasets.
-While XGBoost performs competitively in COADREAD (96.4%), its inconsistency becomes apparent when examining results across datasets, dropping to 71.8% for LGG and 63.5% for STAD.
-DFN shows stronger baseline performance, particularly in STAD (69.1% ROC-AUC), but still falls short of MOTGNN’s 71.2%.
-GCN exhibits moderate performance across datasets and outperforms DFN slightly in LGG and STAD, which suggests some benefit from modeling feature relationships through a graph structure.
-However, GCN constructs only one graph over concatenated multi-omics features, which may weaken omics-specific signals and introduce noise from irrelevant cross-omics interactions.
-In contrast, MOTGNN’s modality-specific graph-based architecture provides additional performance gains, particularly under class imbalance, highlighting its suitability for complex biomedical classification tasks.
-Figure 6.
-F1-score comparison across 20 runs on the imbalanced COADREAD dataset.
-(A) Box plots show median (labeled), interquartile range, and outliers.
-(B) Violin plots show median (labeled) and illustrate score distributions.
-MOTGNN demonstrates both high performance and low variability.
-[…]
-[BoxandViolinPlot]Box and violin plots showing F1-score distributions for different models on COADREAD.
-Figure 5 visually compares model performance across all datasets and evaluation metrics, highlighting MOTGNN’s consistent advantage.
-Although some confidence intervals slightly overlap, MOTGNN consistently achieves higher mean scores.
-The most significant gaps appear in F1-scores for COADREAD, where MOTGNN (87.2%) far exceeds RF (33.4%), with non-overlapping confidence intervals.
-The figure also illustrates that, despite XGBoost’s strong performance in COADREAD (91% accuracy), its accuracy drops considerably in LGG and STAD, underscoring its lack of robustness across datasets.
-Figure 6 highlights model robustness on the imbalanced COADREAD dataset (class ratio 254:78) by comparing F1-scores over 20 independent experiments.
-MOTGNN achieves the highest median F1-score (0.871), outperforming XGBoost (0.794), DFN (0.771), GCN (0.735), and RF (0.349).
-The violin plot reveals a tight distribution around MOTGNN’s median, indicating low variability.
-In contrast, XGBoost, DFN, and GCN show wider interquartile ranges, while RF exhibits low performance and high variance.
-These results underscore MOTGNN’s strength in handling class imbalance, a key challenge in biomedical data analysis.
-Table 6 presents the top 30 most important biomarkers identified by MOTGNN for the COADREAD dataset (Colorectal Adenocarcinoma and Rectal Adenocarcinoma), illustrating the feature-level interpretability of our framework.
-COADREAD encompasses colorectal cancer in a broader context and is commonly analyzed jointly, as colon and rectal adenocarcinomas share overlapping molecular signatures and clinical features.
-This dataset was selected due to its well-characterized molecular landscape and the strong classification performance achieved by our model.
+The reduced feature sizes ($p_{i}^{*}$) vary across omics types due to the nature of XGBoost’s supervised feature selection: the final selected set is the union of features used across all $M$ decision trees, as described in Section 2.2.
+Table 4 summarizes the structural properties of the three constructed feature graphs ($G_{1}$, $G_{2}$, $G_{3}$), where $E_{i}$ denotes the set of edges for Graph $G_{i}$ and the edge-to-node ratio is defined as $m_{i}=\frac{E_{i}}{p_{i}^{*}}$ for $i=1,2,3$.
+Classification Performance.
+Table 5 presents the classification results averaged over 20 independent train-validation-test splits.
+Overall, MOTGNN consistently outperforms all baseline models (XGBoost, RF, DFN, and GCN) across all evaluation metrics (accuracy, ROC-AUC, and F1-score) on all datasets (COADREAD, LGG, and STAD).
+In COADREAD, MOTGNN achieves an accuracy of 93.9%, surpassing XGBoost (91.0%), RF (80.1%), DFN (87.3%), and GCN (87.1%).
+The improvement is even more pronounced in F1-score, where MOTGNN attains 87.2%, marking a 9.3% absolute improvement over XGBoost (77.9%) and a substantial 53.8% advantage over RF (33.4%).
+Given COADREAD’s imbalanced class distribution (254:78), these gains highlight MOTGNN’s robustness under skewed data conditions, where traditional ensemble models degrade substantially.
+The two ablation variants, $\mathrm{MOTGNN}_{\text{gcn}}$ and $\mathrm{MOTGNN}_{\text{rf}}$, both outperform their respective base counterparts (GCN and RF), confirming the benefits of the modality-specific multi-graph design.
+However, the full MOTGNN remains superior, demonstrating that combining XGBoost-based graph construction with the GEDFN module provides complementary advantages for capturing nonlinear and omics-specific dependencies.
+Across all datasets, MOTGNN maintains the highest ROC-AUC values ($\geq$ 96.9% for COADREAD and $\geq$ 3% higher than the best baseline in LGG and STAD).
+While XGBoost performs well in COADREAD (96.4%), its performance declines on LGG (71.8%) and STAD (63.5%), indicating limited generalizability.
+DFN exhibits competitive baseline performance, particularly in STAD (69.1% ROC-AUC), but lacks the ability to capture inter-feature relationships effectively.
+GCN achieves moderate gains by leveraging graph structure, but it constructs a single concatenated graph across omics, which may weaken omics-specific signals and introduce noise from irrelevant cross-omics interactions.
+By contrast, MOTGNN learns individual omics-specific graphs and integrates them through interpretable GNN modules, achieving consistent and robust performance across all datasets.
+These results confirm its effectiveness for complex multi-omics classification tasks.
+Figure 4 compares model performance across all datasets and evaluation metrics, highlighting MOTGNN’s consistent advantage over both baseline and ablation models.
+Although some confidence intervals slightly overlap, MOTGNN consistently attains the highest mean scores across accuracy, ROC-AUC, and F1-score.
+The largest performance gaps appear in F1-scores for COADREAD, where MOTGNN (87.2%) far exceeds RF (33.4%) and outperforms all other methods, including $\mathrm{MOTGNN}{\text{gcn}}$ and $\mathrm{MOTGNN}{\text{rf}}$.
+The figure also shows that, despite XGBoost’s strong performance in COADREAD (91% accuracy), its accuracy drops markedly in LGG and STAD, underscoring limited robustness across datasets.
+Figure 5 further demonstrates model robustness on the imbalanced COADREAD dataset (class ratio 254:78) by comparing F1-scores over 20 independent runs.
+MOTGNN achieves the highest median F1-score (0.871), outperforming XGBoost (0.794), DFN (0.771), GCN (0.735), RF (0.349), and both ablation variants $\mathrm{MOTGNN}{\text{gcn}}$ (0.862) and $\mathrm{MOTGNN}{\text{rf}}$ (0.800).
+The violin plots reveal a compact distribution around MOTGNN’s median, indicating low variability and stable performance.
+In contrast, XGBoost, DFN, and GCN display wider interquartile ranges, while RF shows low performance and high variance.
+The ablation variants outperform their respective baselines but remain below the full MOTGNN, reinforcing that both the GEDFN backbone and XGBoost-based graph generation contribute to the model’s overall strength.
+These results underscore the robustness of MOTGNN under class imbalance, a key challenge in biomedical data analysis.
+Graph Importance and Omics-Level Interpretability.
+Figure 6 displays the relative graph importance for the three constructed graphs.
+The observed edge-to-node ratios (ranging from approximately 2.1 to 2.8, as shown in Table 4) indicate that all graphs are sparse.
+Among them, the miRNA graph ($G_{3}$) shows the highest edge density (2.35-2.79) but the lowest relative importance (24-29%), suggesting that it contributes less to classification performance despite being more connected.
+In COADREAD, the mRNA graph ($G_{2}$) contributes most strongly, with an importance of roughly 42%.
+In LGG and STAD, methylation ($G_{1}$) and mRNA ($G_{2}$) contribute comparably, ranging from 35% to 37%.
+The consistently high methylation importance across datasets (approximately 34-37%) underscores its robustness as a discriminative biomarker, whereas the variability of mRNA importance reflects tissue-specific gene expression dynamics.
+These results demonstrate that MOTGNN effectively identifies the most informative omics modalities while maintaining computational efficiency through sparse graph construction.
+The resulting edge counts remain tractable, supporting scalability to larger or clinically oriented datasets.
+Biomarker Selection and Feature-Level Interpretability.
+Table 6 lists the top 30 most important biomarkers identified by MOTGNN for the COADREAD dataset, illustrating the feature-level interpretability of our framework.
+COADREAD encompasses both colorectal and rectal adenocarcinomas, which are often analyzed jointly because of their overlapping molecular signatures and clinical profiles.
+This dataset was selected for interpretability analysis due to its well-characterized molecular landscape and the strong classification performance achieved by MOTGNN.
 The same feature-ranking procedure can also be applied to the other two cancer types: STAD (Stomach Adenocarcinoma) and LGG (Lower Grade Glioma).
-Among the top-ranked features, Secreted Frizzled-Related Protein 4 (SFRP4) is a well-documented biomarker associated with colorectal cancer (Huang et al., 2010; Yu et al., 2022; Frank et al., 2010; Liu et al., 2020).
-SFRP4 is overexpressed in tumor tissues, and its promoter hypermethylation has been observed as an early event in cancer progression.
-Functionally, it acts as a modulator of Wnt signaling, thereby influencing cell proliferation and tumor growth.
-While our current study focuses on computational identification through feature importance scores, biological validation remains an important next step.
-These selected biomarkers, including SFRP4, serve as potential candidates for further investigation and may contribute to discovering novel diagnostic or therapeutic strategies for diseases.
-Table 6. Top 30 biomarkers identified by MOTGNN for COADREAD by omics type (methylation, mRNA, and miRNA)
+Among the top-ranked features identified for COADREAD, Secreted Frizzled-Related Protein 4 (SFRP4) stands out as a well-documented biomarker associated with colorectal cancer (Huang et al., 2010; Frank et al., 2010; Liu et al., 2020; Yu et al., 2022).
+Multiple studies have reported that SFRP4 is overexpressed in tumor tissues, and its promoter hypermethylation occurs as an early event in cancer progression.
+Functionally, SFRP4 acts as a modulator of Wnt signaling, thereby influencing cell proliferation and tumor growth.
+While the present work focuses on computational discovery through feature-importance analysis, these findings highlight MOTGNN’s potential for guiding downstream biological validation.
+The selected biomarkers serve as promising candidates for further investigation and may ultimately contribute to discovering novel diagnostic or therapeutic strategies for complex diseases.
+Table 6.
+Top 30 biomarkers identified by MOTGNN for the COADREAD dataset (Colorectal and Rectal Adenocarcinoma), grouped by omics type: DNA methylation (meth), gene expression (mRNA), and microRNA expression (miRNA).
+The listed biomarkers represent the biomarkers with the highest feature-importance scores derived from MOTGNN.
 Omics Type Biomarkers
@@
 hsa-mir-206, hsa-mir-30c-2, hsa-mir-425
-Note: The COADREAD dataset refers to Colorectal Adenocarcinoma and Rectal Adenocarcinoma (See data details in Table 2).
-Biomarkers are ranked and selected from three omics types: DNA methylation (meth), gene expression (mRNA), and microRNA expression (miRNA).
-Table 7 reveals the computational trade-offs associated with these performance improvements.
-MOTGNN requires the longest training times, averaging 1.75 minutes for COADREAD and up to 2.39 minutes for LGG, compared to GCN’s 0.59-0.94 minutes, DFN’s 0.3-0.39 minutes, XGBoost’s 0.16-0.25 minutes, and RF’s remarkably fast 0.05 minutes per run.
-However, these computational costs remain practical for real-world applications, especially considering the analysis was conducted using only four CPUs per job.
-The increased runtime is justified by MOTGNN’s superior and stable performance across diverse datasets and metrics.
-Table 7. Computation time comparison across models and datasets (in minutes).
+Table 7 summarizes the computational trade-offs associated with these performance improvements.
+Among all models, MOTGNN requires the longest training time, averaging 1.75 minutes for COADREAD and up to 2.39 minutes for LGG, compared to GCN’s 0.59-0.94 minutes, DFN’s 0.3-0.39 minutes, XGBoost’s 0.16-0.25 minutes, and RF’s remarkably fast 0.05 minutes per run.
+Despite this additional cost, the runtime remains practical for real-world applications, especially considering the analysis was conducted using only four CPUs per job.
+The longer runtime primarily reflects the additional computational steps required for supervised graph construction and hierarchical representation learning.
+These added computations are justified by MOTGNN’s superior and stable performance across diverse datasets and evaluation metrics, demonstrating a favorable balance between efficiency and predictive accuracy.
+Table 7.
+Computation time comparison of MOTGNN and baseline models (XGBoost, RF, DFN, and GCN) across COADREAD, LGG, and STAD datasets.
+Each experiment was conducted using only four CPUs without GPU acceleration, and the reported times are averaged over 20 independent runs.
 Data Model Time (min)
@@
 MOTGNN 1.45
-[…]
 5. Conclusion
@@
 Feature-level importance scores enable the identification of top-ranked biomarkers, which may serve as promising candidates for further biological investigation.
-Meanwhile, omics-level contributions reveal the relative informativeness of each data modality (e.g., methylation, mRNA, or miRNA), offering valuable insight into their role across different disease types.
-Future extensions of the MOTGNN framework may include incorporating additional omics modalities, additional cancer types and non-cancer diseases, or adapting the architecture for more complex prediction tasks, such as multi-class classification or survival analysis.
-Further enhancing interpretability will also be a key direction, especially to better support clinical and biological decision-making.
+Meanwhile, omics-level contributions reveal the relative informativeness of each data modality (e.g., methylation, mRNA, or miRNA), offering valuable insight into their roles across different disease types.
+Together, these interpretability results show that MOTGNN not only enhances prediction performance but also yields biologically meaningful insights that can guide downstream validation and hypothesis generation.
+Future extensions of the MOTGNN framework may include incorporating additional omics modalities, expanding to other cancer types and non-cancer diseases, or adapting the architecture for more complex prediction tasks, such as multi-class classification or survival analysis.
+Further enhancing interpretability and computational scalability will also be important directions, particularly to better support clinical and biological decision-making.
+In the long term, MOTGNN’s modular design and interpretable learning process provide a flexible foundation for advancing integrative multi-omics research and enabling precision medicine applications.
 Data and Code Availability
-The harmonized datasets used in our study and the code for our model are available upon request or publication.
+The original datasets are from TCGA Broad GDAC Firehose (https://gdac.broadinstitute.org/). The harmonized datasets used in our study and the code for our model are available upon publication.
 Acknowledgements.
 This work is partially supported by the National Institute of General Medical Sciences of the National Institutes of Health (NIH/NIGMS) under Award P20GM104420.
-Any findings, comments, and conclusions expressed in this paper do not reflect the views of NIH.